### v3

# Data Augmentation for Time-Series Classification: a Comprehensive Survey

Abstract
Data Augmentation (DA) has emerged as an indispensable strategy in Time Series Classification (TSC), primarily due to its capacity to amplify training samples, thereby bolstering model robustness, diversifying datasets, and curtailing overfitting. However, the current landscape of DA in TSC is plagued with fragmented literature reviews, nebulous methodological taxonomies, inadequate evaluative measures, and a dearth of accessible, user-oriented tools. In light of these challenges, this study embarks on an exhaustive dissection of DA methodologies within the TSC realm. Our initial approach involved an extensive literature review spanning a decade, revealing that contemporary surveys scarcely capture the breadth of advancements in DA for TSC, prompting us to meticulously analyze over 100 scholarly articles to distill more than 60 unique DA techniques. This rigorous analysis precipitated the formulation of a novel taxonomy, purpose-built for the intricacies of DA in TSC, categorizing techniques into five principal echelons: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation. Our taxonomy promises to serve as a robust navigational aid for scholars, offering clarity and direction in method selection. Addressing the conspicuous absence of holistic evaluations for prevalent DA techniques, we executed an all-encompassing empirical assessment, wherein upwards of 15 DA strategies were subjected to scrutiny across 8 UCR time-series datasets, employing ResNet and a multi-faceted evaluation paradigm encompassing Accuracy, Method Ranking, and Residual Analysis, yielding a benchmark accuracy of 88.94 ± 11.83%. Our investigation underscored the inconsistent efficacies of DA techniques, with certain strategies such as Permutation (boasting an average rank of 4.125 and an accuracy pinnacle of 89.20 ± 11.95%, and outperforming in 2 datasets) significantly propelling model performance, whereas others, like Rotation (84.75 ± 12.92%), inadvertently compromised it. Furthermore, we identified that the inherent attributes of datasets exert a profound impact on the success of DA, leading us to provide precise, pragmatic recommendations, derived from our empirical insights, to assist practitioners in aligning their choice of DA techniques with specific dataset nuances.
Index Terms:
Data Augmentation, Time-Series Classification, Machine Learning, Deep Learning.
I Introduction
I-A Background and Motivation
In the realm of machine learning, the quest for robust model performance often hinges on the availability of extensive datasets, a requirement that poses a significant challenge given the frequent scarcity of such data [1]. Data augmentation emerges as a pivotal solution in this context, ingeniously expanding limited datasets by generating additional, analogous data, thereby circumventing the constraints imposed by data insufficiency. This technique, renowned for its cost-effectiveness and efficacy, has gained immense traction across various machine learning disciplines, significantly enhancing model accuracy in environments characterized by limited data. In computational vision, for instance, the augmentation of images is relatively straightforward, with a plethora of simplistic yet potent methods capable of generating discernible images even under conditions of noise addition or partial image cropping [2].
Contrastingly, the augmentation of time series data has not garnered comparable focus. Existing strategies often overlook the intrinsic properties defining time series data, such as temporal correlation, a critical aspect that distinguishes it from other data types. The potential for transforming time series data into frequency or time-frequency domains further complicates the application of conventional augmentation techniques, particularly those borrowed from image or speech processing, leading to potential inaccuracies in synthetic data generation. Moreover, the suitability of data augmentation methods is intricately linked to the specific task, necessitating a nuanced approach that considers the distinct requirements of tasks such as forecasting versus classification within the time series context.
I-B Objectives and Contributions
This manuscript presents an exhaustive exploration into the contemporary landscape of Data Augmentation (DA) for Time Series Classification (TSC), delving into the myriad techniques and methodologies that have sculpted the domain over the preceding decade and furnishing readers with a profound comprehension of its evolution.
• Comprehensive Survey & Novel Taxonomy: Endeavoring to impose order and coherence upon the diverse techniques permeating this field, we unveil a pioneering taxonomy tailored explicitly for DA in TSC. Our classification schema compartmentalizes methods into quintessential categories: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation, thereby facilitating a methodical and streamlined approach to DA for TSC study. This structured paradigm is instrumental in the comparative analysis and judicious selection of techniques, calibrated to the nuances of specific tasks.
• Detailed Evaluation: In a move transcending the conventional survey and taxonomy, this study makes a seminal contribution by undertaking a rigorous evaluation of more than fifteen prominent DA for TSC methodologies. These methods were subjected to a battery of tests employing a proven neural network model, establishing a robust and reliable performance benchmark. Our empirical scrutiny provides an all-encompassing perspective on the relative merits and demerits of these techniques, yielding critical insights of immense value to both the scholarly and practical realms.
• Unified Library: Complementing these efforts, we have orchestrated a Python library, encapsulating the aforementioned time series data augmentation algorithms, designed for effortless implementation. This initiative stems from the recognition of the existing tools’ non-intuitive nature, necessitating a more user-friendly alternative.
The revelations from our investigative journey contribute a panoramic view of the current techniques, simultaneously illuminating prospective avenues for refinement and future inquiry. Our aspiration is to catalyze further innovation in DA for TSC, charting the course for the emergence of increasingly sophisticated, efficient methods adept at navigating the complexities intrinsic to time series data. The principal contributions of this manuscript are encapsulated as follows:
  • We orchestrate a meticulous review of the advancements characterizing the past decade in Data Augmentation for Time-Series Classification (DA for TSC);
  • We advocate a novel, purpose-built taxonomy for DA for TSC, segmenting techniques into quintessential categories: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation;
  • We execute an in-depth evaluation of over fifteen DA for TSC techniques, utilizing a renowned neural network model, and offer an exhaustive discourse on the findings;
  • We introduce a Python compendium of preeminent time series data augmentation algorithms, conceived to present a unified, user-friendly platform for the application of diverse data augmentation strategies to time series datasets.
The rest of this article is organized as follows (Figure 1). In Section 2, we introduce the literature review of DA for TSC. Section 3 presents the survey methodology of our works and discussion of the result. In Section 4, we present the taxonomy for DA techniques, which compartmentalizes the techniques into five principal categories. In Section 5, we introduce the experimental study setup in detail and raise our research questions. After that, the research questions are answered in Section 6 and we discuss the limitations and threats to validity in Section 7. Finally, conclusions are drawn in Section 8.
Figure 1: The Overall Organization of this Survey Paper
II Literature Review
Time Series Classification: The domain of Time Series Classification (TSC) represents a pivotal frontier in data mining, dedicated to the categorization of time series into distinct classes predicated upon historical pattern analysis [3]. Inherent in a time series is a sequence of data points, chronologically ordered and indexed over regular intervals, underscoring the temporal dynamics integral to various practical realms, including health diagnostics, financial market analytics, among others. The intricate nature of time series data, marked by heterogeneous lengths, noise interferences, and nonlinear trajectories, poses substantial analytical challenges.
Data Augmentation: Within the spheres of machine learning and deep learning, data augmentation stands out as a strategy for amplifying the training dataset by generating modified iterations of the original data instances [4]. This process, applicable across diverse tasks such as image recognition, involves transformations such as rotation, scaling, or translation. In the context of TSC, augmentation encompasses techniques like noise injection, slicing, jittering, and scaling, serving dual core objectives: enhancing model generalization to diminish overfitting and addressing class imbalances within datasets.
Despite its established efficacy in various machine learning applications, data augmentation’s integration into TSC remains a vibrant research avenue, necessitating deeper exploration. Historically, data augmentation has proven indispensable in scenarios plagued by data scarcity or imbalance, as evidenced in applications like AlexNet for ImageNet classification [5]. Recent scholarly endeavors have extended these techniques to time series data, albeit with modifications to accommodate the data’s unique characteristics [6, 7, 8, 9].
The singular nature of time series data often precludes the direct application of standard augmentation methods, necessitating adaptations or entirely novel methodologies, particularly for tasks like classification, forecasting, and anomaly detection within time series [10]. The past decade has witnessed a proliferation of TSC algorithms, spurred by the increased availability of time-series data and its ubiquity across tasks necessitating human cognitive processes [11, 12]. The diverse dataset types within the UCR/UEA archive further testify to the expansive applicability of TSC [13].
Addressing data imbalances, a prevalent issue in machine learning tasks, requires strategic data augmentation interventions, such as oversampling or undersampling, to rectify biases towards majority classes and enhance minority class performance. Techniques like SMOTE (Synthetic Minority Over-sampling Technique) generate synthetic examples to balance class distributions, while their counterparts reduce majority class instances [14]. However, these approaches necessitate careful calibration to avoid overfitting or information loss.
Innovative contributions to the field include Guennec et al.’s two-layer convolutional network for TSC [15], Germain et al.’s time series-specific augmentation methods [16], and Pfister et al.’s exploration of wearable sensor data augmentation [6]. Haradal et al. ventured into biosignal classification, employing generative adversarial networks (GAN) for synthetic time series data generation [17]. Fons et al. examined financial time series classification, evaluating various data augmentation strategies with deep learning models [18]. Nonaka et al. proposed ECG-specific data augmentation, enhancing atrial fibrillation classification using single-lead ECG data without modifying the DNN architecture [19]. Goubeaud et al. introduced White Noise Windows, a novel time series data augmentation technique, in 2021 [20].
Despite its advantages, data augmentation is not without its challenges, including the risk of overfitting and the potential ineffectiveness with high-dimensional or complex data structures [4, 21].
II-A Relevant Survey & Study
The scholarly landscape is replete with surveys on data augmentation, spanning domains from image processing [22] to natural language processing [23], though comprehensive studies in the time series sector remain scarce. Iwana et al. conducted an empirical assessment of time series data augmentation methods using neural networks, albeit with a focus limited to classification and excluding frequency domain or externally trained methods [4]. Wen et al. expanded upon this work, incorporating time series forecasting and anomaly detection, though their study utilized a relatively constrained dataset from the Alibaba Cloud monitoring system [21].
While these studies offer valuable insights into data augmentation techniques, there is a conspicuous absence of a holistic view encompassing recent advancements and diverse application fields. This paper seeks to bridge this scholarly gap, providing an exhaustive analysis of the current state of time series data augmentation, and serving as a compass for future research trajectories in this burgeoning field.
III Survey Methodology
This paper adopts a tripartite research methodology encompassing Survey, Evaluation, and Experimental Study Setup. The Survey delineates our systematic approach to literature accumulation and analysis. The Evaluation introduces a novel taxonomy of time series data augmentation, predicated on distinctive characteristics. The Experimental Study Setup discusses the authoritative datasets employed, evaluation metrics, and potential variables influencing outcomes, thereby ensuring a rigorous analysis of time series data augmentation methodologies.
III-A Paper Collection Criteria of the Survey
The literature aggregation for this study adheres to stringent inclusion criteria, encompassing:
  • Papers elucidating the fundamental concepts of data augmentation for time-series classification or its associated facets.
  • Studies proposing enhancements to existing data augmentation methodologies, tools, or frameworks within the realm of time-series classification.
  • Contributions introducing benchmarks or datasets tailored for data augmentation in time-series classification.
  • Literature offering comprehensive reviews or surveys on the contemporary landscape of data augmentation for time-series classification.
Exclusions pertain to studies applying relevant techniques outside the time-series context and those focusing on time-series classification, forecasting, or anomaly detection without integrating data augmentation methodologies.
III-B Paper Collection Methodology of the Survey
Our literature retrieval process commenced with exhaustive searches across renowned scientific repositories like Google Scholar and arXiv, utilizing precise keyword combinations. Recognizing the interdisciplinary application of time series data augmentation, we adopted a snowballing strategy for comprehensive coverage [24]. The accumulated literature was subsequently categorized as Survey/Review papers, Research papers, or Methods papers, following established academic conventions [25, 26, 27].
The “keywords” used for searching and collecting are listed below. [DA for TSC] means Data Augmentation for Time-Series Classification, which is the focus of this research project. The search strategy encompassed several iterations, as detailed below:
  • DA for TSC + Survey / Review
  • DA for TSC + Techniques / Methods
  • DA for TSC + Evaluations / Comparison
III-C Result and Discussion of the Survey
TABLE I: Paper Query Results
Paper query results Key Words Hits Title
[DA for TSC] + Survey / Review 20 15 3
[DA for TSC] + Techniques / Methods 258 92 57
[DA for TSC] + Evaluations / Comparison 70 35 16
Query - - 76
Snowball - - 26
Overall - - 102
The terms “Hits” and “Title” refer to the number of papers returned by the search, as well as the number of papers that remain after title and body framing, respectively. The search strategy yielded 102 relevant publications, with a significant proportion originating from keyword searches and the remainder from snowballing. The annual publication trends indicate a burgeoning interest in this research area, particularly post-2019.
Figure 2: Annual Publication Trends in Data Augmentation for Time-Series Classification
The distribution of publications across various publishers underscores the dominance of IEEE and arXiv, with significant contributions also emanating from Elsevier, MDPI, ACM, and Springer.
Figure 3: Publisher Distribution of Publications in Data Augmentation for Time-Series Classification
A geographical analysis of the publications reveals a preponderance of contributions from the United States, followed by China, Japan, Germany, and the United Kingdom, with disparate representation from other global regions.
Figure 4: Publication Count by Country in Data Augmentation for Time-Series Classification
An industry-centric examination, based on the North American Industry Classification System (NAICS), highlights a predominant application within the Information Technology sector, followed by Healthcare, Agriculture, Astronomy, Finance, and Construction, indicating the methodologies’ broad applicability.
Figure 5: Publication Count by Industry in Data Augmentation for Time-Series Classification
IV Data Augmentation for Time-Series Classification Techniques
The taxonomy presented in this study for data augmentation techniques applicable to time-series classification (TSC) emerges from an exhaustive literature survey and rigorous research. It compartmentalizes the techniques into five principal categories: Transformation-Based Methods, Pattern-Based Methods, Generative Methods, Decomposition-Based Methods, and Automated Data Augmentation Methods. This categorization, meticulously refined and expanded upon from preceding studies [10, 4, 21], lays the groundwork for our in-depth exploration and experimentation.
Several factors underpin the formulation of this taxonomy. Primarily, it encompasses a broad spectrum of contemporary methods, thereby paving the way for domain-centric experimental undertakings in specialized segments like the time and frequency domains. Moreover, it offers an organized synopsis of the myriad techniques used for augmenting time-series data, thereby simplifying the comparative analysis and comprehension of diverse methods. Such a structured approach is indispensable for researchers and practitioners alike, guiding them in the judicious selection of data augmentation techniques tailored to their specific requirements. Furthermore, by highlighting the existing lacunae and prospects in the current scenario, it spurs the innovation of novel techniques.
The distinct categories within this taxonomy epitomize different methodologies in data augmentation. Transformation-Based Methods involve the alteration of original data via various transformations. Pattern-Based Methods, on the other hand, generate new instances through the extraction and subsequent recombination of patterns inherent in the original data. Generative Methods utilize generative models to synthesize new data instances. Decomposition-Based Methods, in contrast, entail the disintegration of the original time-series into multiple components, which are then individually augmented and reassembled. Automated Data Augmentation Methods streamline the augmentation process, often leveraging data-driven insights to ascertain the most efficacious augmentation strategy.
The ensuing subsections provide an exhaustive discourse on each category, elucidating their respective methodologies, merits, and potential use-cases. This detailed analysis serves dual purposes: guiding users in the selection of appropriate data augmentation strategies and fostering the conception of avant-garde methods in forthcoming research endeavors.
{forest}
[…]
for Time-Series
Classification, align=center, before computing xy=l=2cm [Transformation-Based
Methods, align=center [Magnitude [Jittering [6, 28, 29, 30, 31, 32, 33]] [Rotation [6, 34, 32, 33]] [Scaling [6, 7, 35, 29, 30, 32]] [Magnitude Warping [6, 35, 36, 30]] [White Noise Windows [20]] ] [Time [Permutation [6, 34, 30]] [Slicing [15, 29]] [Time Warping [35, 32]] [Window Warping [15, 37, 7]] [Window Slicing [15]] ] [Frequency [Frequency Warping [VTLP [38]] [SFCC [39]] ] [Frequency Slicing [29]] [Random Noise Boxes [40]] ] [Time-Frequency [Frequency Transform [STFT [41]] ] [Surrogate Data Augmentation [42]] ] ] [Pattern-Based
Methods, align=center [Magnitude [TS-DENSE [43]] [RIM [44]] [SPAWNER [45, 46]] [ECGAug [47]] ] [Time [Guided Warping [48] [RGW] [RGWs] [DGW] [DGWs] ] [Interpolation [Interpolation [49]] [Shapelets-based SMOTE [14]] [DTW-SMOTE [50]] ] [DBA [16, 51]] [ASP [52]] [Mapping Label [53]] [StyleTime [54]] [DTW-MERGE [55]] ] [Frequency [EMDA [56]] [SFM [57]] ] [Time-Frequency [Sample Combining [58]] [Audio Augmentation [59, 60]] ] ] [Generative
Methods, align=center [Statistical Models [Gaussian trees [61]] [GRATIS [62]] ] [Deep Learning, l=2cm [GAN [GAN [17, 63, 64, 36, 65, 66]] [Wasserstein GAN [67]] [Conditional GAN [68, 69]] [EmotionalGAN [70]] [BiLstM-CNN GAN [71]] [Time Series GAN [72, 73, 74, 75, 76]] [GRU-CNN GAN [77]] [Imputation Balanced GAN [78]] [PlethAugment [79]] ] [Encoder Decoder [AE [80, 81]] [VAE [82, 83, 84, 85, 86, 87]] ] [ODENet [88]] ] [Dynamical System [89, 90]] [Signal Augmentation [91]] [IMDA [92]] ] [Decomposition-Based
Methods, align=center [EMD [93, 94]] [STL [95, 96, 97, 98]] ] [Automated Data
Augmentation
Methods, align=center [RandECG [99]] [Fast AutoAugment [82]] [Multiple Approaches [100]] [Hybrid Framework [101]] [JobDA [102]] [Adaptive Weighting Scheme [103]] [MODALS [104]] ] ]
Figure 6: Taxonomy of Techniques of Data Augmentation for Time-Series Classification
IV-A Transformation-Based Methods
Transformation-based methods hinge on the principle of applying an array of transformations to the original time-series data. The objective is to retain the fundamental structure while introducing nuanced variations, thereby bolstering model resilience. These methods, characterized by their simplicity and intuitiveness, typically encompass elementary mathematical operations, making them versatile across diverse data types and domains. The scope of transformations extends to various data facets, culminating in four distinct sub-categories.
Figure 7: Illustrative Examples of Transformation-Based Methods on the ECG5000 Dataset (Generated instances are depicted in orange)
IV-A1 Magnitude Transformations
Magnitude transformations alter the amplitude of time-series data, an approach especially pertinent in contexts where the data’s magnitude is informative, such as signal processing or audio analysis.
For instance, the study by Terry et al. [6] delves into data augmentation techniques like Jittering, Scaling, Rotation, Magnitude-Warping, etc, applied for Parkinson’s disease movement classification using wearable sensor data. Research underscores that an amalgamation of DA methods markedly propels classification efficacy, jumping from a baseline of 77.52% to 86.88%. This finding signals a promising trajectory for wearable sensor data augmentation. Similarly, Shuang et al. [35] harnesses Scaling, Magnitude Warping, and other techniques on I/Q signal data, culminating in a peak recognition accuracy of 93.68% and successful differentiation between QAM16 and QAM64 signals. Additionally, Goubeaud et al. [20] introduce White Noise Windows (WNW), a novel augmentation technique that fortifies classification methods and mitigates overfitting.
IV-A2 Time Transformations
Time transformations adjust the temporal elements of the data, such as data point sequencing or timing. They are instrumental when the data’s temporal dynamics are crucial, as seen in tasks like time-series classification, forecasting, or sequence prediction.
In addition to the aforementioned magnitude transformations, the work of Terry et al. [6] also integrates time-based data augmentation techniques such as Permutation and Time-Warping, further demonstrating their utility in enhancing Parkinson’s disease movement classification through wearable sensor data. Similarly, the study by Shuang et al. [35] employs Time-Warping among its suite of techniques, showcasing its effectiveness in improving the recognition accuracy of I/Q signal data. These instances underline the broad applicability and potential of time transformations in augmenting data for varied classification tasks.
In the realm of construction equipment IMU data, Rashid et al. [7] innovates with Window-Warping (WW), a technique that synthesizes training data, thereby enhancing machine learning model proficiency in activity recognition tasks, with K-Nearest Neighbors (KNN) witnessing substantial improvements. Furthermore, Guennec et al. [15] discusses two augmentation methods for time series classification with Convolutional Neural Networks: Window Slicing (WS) and Window Warping (WW). These strategies, coupled with a semi-supervised Dataset Mixing technique, bolster classification performance, particularly in smaller datasets.
IV-A3 Frequency Transformations
Frequency transformations tinker with the time-series data’s frequency components, a strategy quintessential in scenarios where the data’s frequency content is vital, such as in spectral or vibration analysis.
In 2023, Yang et al. [39] pioneers Stratified Fourier Coefficients Combination (SFCC), a method that merges original Fourier coefficients to augment time series datasets, thereby boosting deep neural network performance. This efficient technique achieves state-of-the-art results. In a similar vein, Goubeaud et al. [40] unveils Random Noise Boxes (RNB), an innovative augmentation technique optimized for spectrogram classification. While particularly potent for ECG and sensor datasets, its efficacy diminishes for motion datasets, underscoring the need for vigilant application to avert performance deterioration.
IV-A4 Time-Frequency Transformations
Time-frequency transformations concurrently tweak both temporal and frequency aspects. This dual approach is advantageous when both elements are pivotal, enabling the generation of new instances that preserve the original data’s temporal and frequency structures.
In 2019, Lee et al. [42] employs data augmentation techniques, notably Amplitude Adjusted Fourier Transform (AAFT) and Iterative Amplitude Adjusted Fourier Transform (IAAFT), to concoct surrogate data conducive to deep learning. These strategies maintain the time-domain statistics and power spectrum of the original time series, thereby enriching the training of pre-trained Convolutional Neural Networks (CNNs). The research corroborates the efficacy of these methods in automated rehabilitation classification tasks, suggesting avenues for further exploration.
IV-B Pattern-Based Methods
Pattern-based methods concentrate on discerning and capitalizing on specific patterns or structures inherent in the time-series data. These patterns, whether local or global, simple or complex, often involve machine learning models for identification and utilization. The hallmark of this category is its emphasis on the data’s intrinsic patterns, a focus that proves invaluable in contexts with pronounced structural or pattern-based components, such as electrocardiograms (ECG) or stock market data.
Figure 8: Illustrative Examples of Pattern-Based Methods on the Strawberry Dataset (Generated instances are depicted in orange)
IV-B1 Magnitude Pattern Mining
This niche capitalizes on significant local structures or motifs within the data, exploiting these elements to spawn new, realistic data instances.
In 2019, Kamycki et al. [45] introduces SPAWNER, an innovative data augmentation method that aligns input sequences and restricts the warping path to produce synthetic examples, thereby enhancing the performance of time-series classifiers on benchmark datasets. In a parallel study, Aboussalah et al. [44] employs Recurrent In-fill Masking (RIM) in 2022, a novel augmentation method demonstrated to bolster time series classification task performance across synthetic and real-world datasets, outperforming traditional training and other augmentation strategies.
IV-B2 Time Pattern Mining
These methods adjust the data’s temporal characteristics, maintaining the original data’s temporal structure while generating new instances.
Forestier et al. [16] advocates for a data augmentation method that utilizes Dynamic Time Warping (DTW) to synthesize examples by averaging time series sets, a method especially potent in scenarios with limited example availability. The study confirms that augmenting full datasets typically enhances accuracy, with notable improvements in 56 of the 85 tested datasets. Similarly, Iwana et al. [48] and Cheolhwan et al. [49] introduce novel methods that leverage time warping and cubic interpolation, respectively, to generate additional training samples effectively, thereby optimizing model performance across various datasets.
IV-B3 Frequency Pattern Mining
Frequency Pattern Mining methods focus on manipulating the frequency components of time-series data. These methods are pivotal when the data’s frequency content is crucial, enabling the generation of new instances that preserve the original data’s frequency structure.
In [56], Loris et al. introduce a technique known as EMDA, or Equalized Mixture Data Augmentation, a method premised on spectrograms. EMDA operates by forging a new spectrogram through the weighted average of two randomly selected spectrograms from the same class, incorporating perturbations like time delay and equalizer functions. This strategy, as per the authors, enhances the diversity and robustness of audio signals, thereby fortifying classification performance. Conversely, Cui et al. [57] advocates for Stochastic Feature Mapping (SFM), a data augmentation method tailored for deep neural network acoustic modeling. SFM, a label-preserving transformation, statistically transmutes one speaker’s speech data into another’s within a specified feature space. The essence of SFM is to amplify pattern variations via transformations, thereby bolstering the neural networks’ classification invariance and generalization capabilities.
IV-B4 Time-Frequency Pattern Mining
Time-Frequency Pattern Mining methods adjust both time and frequency components of the data. This dual-pronged approach is beneficial when both the data’s temporal dynamics and frequency content are integral, as seen in domains like speech recognition or music analysis.
In 2020, Zhao et al. [58] proposes a unique method that transitions time-domain data into the frequency domain via a discrete cosine transform (DCT), subsequently generating artificial data by merging different frequency bands. These synthesized signals are then reverted to the time domain through an inverse DCT, with experimental outcomes underscoring the method’s efficacy. In a related study, Nanni et al. [59] employs data augmentation strategies to enhance the classification accuracy of animal sounds using convolutional neural networks (CNNs). These strategies, encompassing transformations of audio signals and their spectrograms, utilize techniques like pitch shift, noise addition, and equalized mixture. The findings affirm that data augmentation bolsters the accuracy and generalization capabilities of CNNs, with a fusion of different CNNs yielding optimal results.
IV-C Generative Methods
Generative Methods constitute a separate class within the taxonomy of time-series data augmentation techniques. These methods leverage generative models to craft new time-series instances, diverging from transformation-based and pattern-based methods by incorporating a learning phase. During this phase, the model discerns the data’s underlying distribution, subsequently spawning new instances from this distribution. This category is invaluable when the objective is to generate diverse, realistic synthetic instances that mirror the original data’s statistical attributes.
Figure 9: Illustrative Examples of Generative Methods (GAN) on the FordB Dataset (Generated instances are depicted in orange)
IV-C1 Statistical Models
Statistical Models utilize generative models with pre-established structures to produce new time-series instances. These models are advantageous when the data’s underlying distribution is known or can be reasonably inferred, facilitating the generation of new instances that adhere to the original data’s distribution.
In [61], a Parsimonious Mixture of Gaussian Trees (2MoGT) model is presented by Cao et al., designed for data augmentation in scenarios with imbalanced and multimodal time-series classification. This model employs Chow-Liu trees to encapsulate dependencies among time-series features and a Gaussian mixture model to accommodate multimodality. It surpasses alternative oversampling methods, especially in datasets with multimodal minority class distributions, and minimizes storage complexity. Additionally, Kang et al. [62] introduces GRATIS, a data augmentation technique that generates time series with adjustable characteristics. Utilizing a mixture autoregressive (MAR) model, attuned to specific features, it spawns diverse time series. The resulting series exhibit diversity, attributable to the mixtures’ adaptability and random parameter configurations, proving especially beneficial for analyzing collections of time series with varying lengths, scales, and characteristics.
IV-C2 Deep Learning Models
Deep Learning Models employ generative models without predefined forms to create new time-series instances. These models are beneficial when dealing with data containing complex, non-linear, or multi-dimensional patterns, as they can generate new instances that encapsulate these intricate patterns.
In [17], a data augmentation strategy based on Generative Adversarial Networks (GANs) is presented by Haradal et al. for biosignal classification. This method incorporates a Recurrent Neural Network (RNN) with LSTM units for the generation of time-series data. Its efficacy is validated by an uptick in classification accuracy, underscoring its viability for synthesizing biosignals and enhancing biosignal classification. Conversely, in 2022, Garcia et al. [64] applies GANs for data augmentation, synthesizing light curves from variable stars. This methodology, merged with an innovative resampling technique and evaluation metric, markedly elevates the classification accuracy of variable stars, proving particularly potent for unbalanced datasets and in pinpointing GAN-overfitting scenarios.
Several researchers have also harnessed specialized GANs for data augmentation. For instance, Ehrhart et al. [69] employed a conditional Generative Adversarial Network (cGAN) with a diversity term for data augmentation, utilizing an LSTM for the generator and an FCN for the discriminator. This method synthesized moments of stress from physiological sensor data, bolstering classifier performance and emerging as a promising technique for stress detection in wearable physiological sensor data. Moreover, in 2023, Yang et al. [72] introduced a time-series Generative Adversarial Network (TS-GAN) for augmenting sensor-based health data, generating synthetic time-series data that preserves the original data’s temporal dependencies, thereby enhancing model performance. The study asserts that TS-GAN surpasses conventional augmentation methods in maintaining data consistency. Additionally, Deng et al. [78] implemented a novel data augmentation strategy, IB-GAN, merging imputation and balancing techniques. It boosts classification performance by synthesizing synthetic samples for under-represented classes, averting the mode collapse often seen in standard GANs. This approach records substantial performance enhancements against cutting-edge baselines. In [70], an emotion state identification framework and ECG signal augmentation algorithm based on GAN are proposed by Chen et al. EmotionalGAN is utilized to synthesize synthetic ECG data, which is then amalgamated with the original training set, enhancing emotion state classification. The findings indicate a notable improvement in classification performance with the integration of EmotionalGAN-synthesized data.
Models based on the Encoder Decoder structure have also gained traction in Data Augmentation for Time-Series Classification. Moreno et al. [85] employs Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) for data augmentation, enhancing prediction accuracy on small datasets, with substantial gains realized through computational effort. The study also introduces alterations to balance class representation in synthetic samples. Zha et al. [80] unveils a straightforward yet potent self-supervised model for time series generation, termed the Masked Autoencoder with Extrapolator (ExtraMAE). The synthetic time series crafted by ExtraMAE exhibit remarkable practicality and fidelity, eclipsing all benchmarks while remaining more compact and expedient in a minimalist design. Furthermore, Hsu et al. [87] introduces two Variational Autoencoder (VAE)-based data augmentation strategies for unsupervised domain adaptation in Automatic Speech Recognition (ASR). These methods, encompassing nuisance attribute replacement and soft latent nuisance subspace perturbation, alter nuisance attributes of speech, achieving approximately a 35% absolute Word Error Rate (WER) reduction.
IV-D Decomposition-Based Methods
Decomposition-Based Methods disassemble the original time-series into multiple components, augment each component independently, and subsequently reassemble them to produce new instances. This unique tactic of deconstructing the time-series data and individually addressing each component proves advantageous when various components of the data necessitate distinct augmentation techniques.
In [93], a data augmentation strategy predicated on Empirical Mode Decomposition (EMD) and an Integrated LSTM neural network is showcased by Li et al. Raw hand motor signals are fragmented by EMD, followed by the augmentation of the decomposed signals through LSTM. The method, trialed on the Ninaweb public dataset, demonstrated a 5.2% ascension in classification accuracy compared to benchmark outcomes. In 2019, Wen et al. [96] introduces RobustSTL, a pioneering algorithm for time-series decomposition that tackles anomalies, sudden trend shifts, seasonality alterations, and noise. It employs bilateral filtering for noise eradication, LAD regression for trend extraction, and non-local seasonal filtering for seasonality determination. The algorithm surpasses existing solutions in managing complex time-series data. An extension of RobustSTL is presented in [98] by Wen et al., designed to accommodate multiple seasonalities in time series data, utilizing a specialized generalized ADMM algorithm. This method outperforms leading algorithms like MSTL, STR, TBATS in terms of effectiveness and computational efficiency, with future explorations geared towards multi-resolution seasonal-trend decomposition for further enhancements.
IV-E Automated Data Augmentation Methods
Automated Data Augmentation Methods strive to automate data augmentation processes, often discerning the optimal augmentation strategy directly from the data. This category distinguishes itself with its focus on automation and adaptability, ideal for circumstances where manual selection and adjustment of augmentation techniques are either impractical or not optimal.
In 2022, Park et al. [82] harnesses PoseVAE and Fast AutoAugment for data augmentation, with PoseVAE rejuvenating skeleton data based on initial human poses, thereby generating stable and novel data. Fast AutoAugment fine-tunes the augmentation policy employing RNN-based reinforcement learning. The research concludes that these methodologies can effectively augment skeleton data, boosting model performance.
[…]
It fine-tunes four universal data transformation operations to conform to various data modalities, creating challenging examples and employing additional loss terms to refine label-preserving transformations. It manifests effectiveness across diverse modalities, including text, tabular, time-series, and images.
V Experimental Study Setup
In this section, we detail the comprehensive comparative assessments conducted, utilizing various data augmentation techniques. These evaluations are designed to provide an empirical comparison between different data augmentation strategies. The primary focus of the comparison is to assess the influence of these methods on the performance of neural network models, especially in the context of time series datasets. This approach allows for a deeper understanding of how these techniques can be optimized to enhance model performance, offering a broader view of their potential impact across various applications.
V-A Datasets
This research encompasses the application of several time-series classification techniques to diverse datasets from the University of California, Riverside (UCR) time-series classification repository [13]. The UCR repository is a rich collection of time-series datasets, extensively utilized within the academic community for benchmarking time series classification algorithms. For this study, eight representative categories were chosen, each representing a unique type of time-series data:
• Image (Yoga): Time-series representations of images, posing challenges in capturing spatial-temporal dynamics.
• Sensor (FordB): Sensor data, often fraught with noise, embodying the challenges of real-world data collection environments.
• Simulated (CBF): Simulated datasets, offering controlled environments where specific time-series phenomena can be isolated and studied.
• Device (ScreenType): Illustrating the challenges posed by device-specific readings, which may vary based on individual device characteristics.
• Motion (GunPointAgeSpan): Encompassing the complexities of motion data and its unique temporal dynamics.
• Spectro (Strawberry): Spectrograms and related time-frequency representations, posing challenges due to their multi-modal data nature.
• ECG (ECG5000, ECG200): Medical time-series data, critical for its implications in health and diagnostics, yet complex due to its sensitivity to minor temporal shifts.
TABLE II: Selected datasets from the 2018 UCR Time Series Classification Archive
Dataset Type Train Test Class Length
CBF Simulated 30 900 3 128
ECG200 ECG 100 100 2 96
ECG5000 ECG 500 4500 5 140
FordB Sensor 3636 810 2 500
GunPointAgeSpan Motion 135 316 2 150
ScreenType Device 375 375 3 720
Strawberry Spectro 613 370 2 235
Yoga Image 300 3000 2 426
The datasets selected for this study underwent preprocessing to ensure compatibility with the data augmentation techniques. In Table II, ”Class” refers to the number of distinct categories or labels within the dataset. It reflects the granularity and diversity of events or phenomena that the dataset captures. A higher number of classes generally suggests a more complex classification task. ”Length” denotes the duration of each individual time series in the dataset, representing the number of time-points or observations for each sequence. This factor is crucial as longer sequences might capture more detailed dynamics but also introduce challenges in computational complexity and potential for overfitting.
Firstly, all datasets were rescaled to ensure that the minimum value in the training set is -1 and the maximum is 1. This normalization step helps preserve relative values and maintain consistency across different datasets. Additionally, to accommodate batch training procedures requiring input data of identical dimensions, datasets of varying lengths were zero-padded. This process involved adding zeros at the end of shorter sequences to match the length of the longest sequence in the dataset. Lastly, we addressed missing values, if present, by substituting them with zero, a standard approach that maintains the dimensional integrity of the data while ensuring that missing values do not influence the learning process.
V-B Evaluated Network Models
For our evaluation, we employed the deep Convolutional Neural Network (CNN) model, ResNet, known for its utilization of residual connections. The hyperparameters for the 1D ResNet used in this evaluation were informed by research conducted by Fawaz et al. [3] and Iwana et al. [4], where they noted improvements in time series classification using data augmentation.
As shown in Figure 10, our implementation of ResNet diverges from the original model [105] by incorporating only three residual blocks with varying filter lengths and omitting max pooling. The first block employs three layers of 64 1D convolutions, the second block contains three layers of 128 1D convolutions, and the third block features three layers of 128 convolutions. Each residual block has three convolutional layers with filter sizes of 8, 5, and 3, respectively. Batch normalization [106] and ReLU activation functions follow each convolution. Additionally, the residual connections merge the input of each residual block with the input of the subsequent block through an addition operation. The final two layers of ResNet comprise a Global Average Pooling (GAP) layer and an output layer with softmax activation.
All network parameters underwent initialization using Glorot’s Uniform method [107]. We trained the network using the Adam optimizer, with an initial learning rate of 0.001 and exponential decay rates for the first and second momentum estimates at 0.9 and 0.999, respectively, following Fawaz et al.’s methodology [3].
Figure 10: The architectures of the neural networks used to evaluate the data augmentation methods [4]
To maintain consistency in the comparison of data augmentation methods, we ensured each model underwent the same 10,000 iterations, reducing the learning rate by 0.1 on plateaus of 500 iterations. The number of patterns in the training dataset was quadrupled compared to the original dataset. The choice of 10,000 iterations stemmed from observations indicating that the training loss for the model on each dataset typically converged before this point.
Moreover, as the number of iterations remained constant, the control experiment without augmentation effectively repeated the same patterns to match the augmented experiments. Regarding hyperparameters selection, we adhered to the best practices documented in literature as closely as possible, avoiding excessive adjustments to maintain a fair comparison of the data augmentation effects.
V-C Evaluated data augmentation methods
TABLE III: Evaluated Data Augmentation Methods and Their Categories
No. Data Augmentation Method Category
1 None -
2 Jittering [6] Transformation-Based
3 Rotation [6] Transformation-Based
4 Scaling [6] Transformation-Based
5 Magnitude Warping [6] Transformation-Based
6 Permutation [6] Transformation-Based
7 Random Permutation [6] Transformation-Based
8 Time Warping [6] Transformation-Based
9 Window Slicing [15] Transformation-Based
10 Window Warping [15] Transformation-Based
11 SPAWNER [45] Pattern-Based
12 wDBA [16] Pattern-Based
13 RGW [48] Pattern-Based
14 RGWs [48] Pattern-Based
15 DGW [48] Pattern-Based
16 DGWs [48] Pattern-Based
17 GAN [66] Generative
The data augmentation methods scrutinized in this study are delineated in Table III. The criterion for their selection was predominantly their recurrent citation in scholarly literature and the accessibility of their open-source implementations, which facilitates reproducibility and comparative scrutiny. Amidst a plethora of techniques, some entailing elaborate combinations and complexities, the endeavor was to encapsulate those epitomizing foundational and seminal contributions in the field. This strategic choice ensures the study’s resonance, relevance, and alignment with contemporary scholarly discourse and benchmarks.
Each method under evaluation was configured using parameters from their respective seminal papers, thereby preserving the authenticity of their original design and ensuring an equitable evaluation landscape. These methods were appraised based on their efficacy in enhancing the performance of the ResNet model across diverse time series datasets.
V-D Evaluation Metrics
In addressing the intricate task of evaluating algorithm reliability within the domain of time series analysis, we acknowledge the absence of a singular metric universally applicable across all conceivable applications. This context underpins our selection of evaluation metrics, aiming to bridge the gap between general applicability and the specific demands of our study. This research used a multifaceted evaluation metric system to assess the effectiveness of data augmentation methods. Each metric contributes a unique evaluative perspective, cumulatively offering a holistic insight into each method’s efficacy. The metrics incorporated are as follows:
1. Accuracy (Acc): Accuracy, an intuitive and globally recognized metric, gauges the ratio of correctly predicted instances in all prognostications. We have opted for Acc as a primary metric due to its relevance in scenarios where the correct identification of both true positives and true negatives is of paramount importance. This choice is informed by the nature of our balanced dataset, where equal importance is placed on each class and the distinct types of classification errors are deemed equally significant. In scenarios where false negatives and false positives are critically important, alternative metrics such as the F1-score might be preferred; however, given the balanced nature of our datasets and the equivalence of misclassification impacts, Acc was deemed the most appropriate metric for our analysis of data augmentation in time-series classification tasks.
2. Method Ranking: To facilitate a more nuanced comparative analysis of the data augmentation methods under investigation, we have adopted a ranking paradigm based on accuracy scores. This approach not only simplifies the identification of the most effective strategies but also establishes a clear hierarchy of performance across different datasets. A lower ranking is indicative of superior methodological performance, offering a structured framework to assess the relative success of each data augmentation strategy.
3. Residual Analysis: Residual plots were instrumental in juxtaposing the performance of each method with the baseline (None). Herein, a residual denotes the deviation between the actual and forecast accuracy. Positive residuals signify outperformance over the baseline, whereas negative residuals indicate underperformance. These plots provide a visual conduit for assessing the relative performance increments or decrements attributable to each data augmentation strategy.
This multimetric evaluation approach underpins the study’s commitment to a nuanced, rigorous, and exhaustive assessment, fortifying the conclusions regarding the effectiveness of each data augmentation strategy.
V-E Research Questions and Objectives
Navigating the complexities of time-series data augmentation necessitates a profound understanding of the subtleties and ramifications of various methods. This study is predicated on the pursuit of clarity in this domain, aiming to elucidate the efficacy and implications of diverse data augmentation strategies through the following research questions:
V-E1 RQ1: What is the impact of data augmentation on model performance?
This pivotal question seeks to explore the influence of data augmentation on model efficacy, probing the disparities in outcomes engendered by different techniques. It bifurcates into more specific inquiries:
V-E1a RQ1.1: What is the extent of the universality of data augmentation strategies?
This inquiry is dedicated to discerning the consistency of data augmentation’s effectiveness across various contexts, highlighting any discrepancies and underlying rationales for performance enhancement or degradation.
V-E1b RQ1.2: Which data augmentation strategy is most efficacious for TSC?
This probe focuses on identifying the preeminent strategy among the evaluated techniques, offering critical insights for practitioners in the time-series domain.
V-E2 RQ2: How do intrinsic dataset characteristics influence the effectiveness of data augmentation and subsequent recommendations?
This exploration delves into the nuanced interplay between dataset attributes and augmentation efficacy, aiming to extrapolate actionable insights for optimal method application based on specific dataset traits. The objective is to equip practitioners with informed, tailored augmentation strategies, thereby maximizing model performance and circumventing the pitfalls of indiscriminate augmentation application.
The ensuing chapter, “Results Analysis” will systematically address these queries, contributing to a robust, comprehensive understanding of time-series data augmentation and its diverse implications in machine learning.
VI Results Analysis
In this section, we undertake a methodical examination of the research questions posited at the outset. Our first point of inquiry assesses the universality and optimality of diverse data augmentation techniques (RQ1). Subsequently, we explore the nuanced interdependencies between dataset attributes and the efficacy of these augmentation methods, providing customized guidance for practitioners (RQ2). This discourse contributes a lucid, empirically grounded viewpoint on the dynamic between data augmentation strategies and time-series classification.
VI-A Answer to RQ1.1: How universal are data augmentation techniques?
TABLE IV: Performance of Different Data Augmentation Techniques
Aug Result
[…]
Aug Result
[…]
Aug Result
[…]
Figure 11: Performance of Different Data Augmentation Techniques
[…]
VI-B Answer to RQ1.2: Which data augmentation technique is optimal for TSC?
Figure 12: Compare the average ranking of Data Augmentation Technique
Figure 13: Accuracy Difference (Compared to None) for each Data Augmentation Technique
VI-B1 Transformation-Based Methods
Transformation-Based methods concentrate on subtle modifications to the original data’s structure, avoiding the creation of entirely new instances. Techniques encompassing Scaling, Warping, and Rotation are categorized within this domain.
Referring to Figure 12, it is noteworthy that six out of the top ten methods are from the Transformation-Based family, suggesting their superior applicability in time-series classification tasks. These methods excel in enhancing data variability while meticulously maintaining the fundamental characteristics and sequences inherent in the original data. For example, both Permutation and Random Permutation diversify the dataset by rearranging time series segments, avoiding extreme modifications that could distort the data’s essential structure. Similarly, Magnitude Warping and Scaling perform elastic transformations, ensuring the fundamental data features remain unaltered. The success of these strategies reaffirms the criticality of preserving the innate temporal dynamics in time-series data while introducing nuanced variations.
VI-B2 Pattern-Based Methods
Pattern-Based methods focus on extracting or amalgamating patterns from the original data, generating new instances that may introduce novel structural elements.
Despite their potential, Pattern-Based methods demonstrated a relatively subdued performance in our evaluation. For example, the SPAWNER technique, despite its innovative approach of merging patterns and infusing noise, may inadvertently introduce a level of convolution to the data, potentially detracting from its original structure. The synthesis of patterns, while creative, risks generating data points that significantly deviate from the original data configurations. The efficacy of these methods may depend on specific dataset characteristics, such as inherent correlations or unique time-series motifs, which could be disrupted unintentionally.
VI-B3 Generative Methods
Generative methods, including models like GANs, are employed to generate new samples that, while potentially divergent structurally from the original data, are designed to be statistically consistent.
In our analysis, GAN, representing Generative Methods, achieved a middle-ground ranking. Despite GANs’ capacity to generate diverse datasets, the intricacies of time-series data present a formidable challenge, particularly in maintaining temporal coherence while fostering variations. However, their ability to generate new data could be particularly beneficial in scenarios with limited data, highlighting their relevance despite a moderate ranking.
VI-B4 Detailed Analysis of Top-Performing and Underperforming Data Augmentation Techniques
Our findings, as depicted in Figure 12 and Figure 13, reveal that Permutation, Random Permutation, and RGWs surpass our benchmark performance. Their consistent presence at the top is not solely a testament to their effectiveness but also possibly indicative of their inherent capability to maintain, and potentially amplify, crucial data features during augmentation. These methods seem to introduce coherent variations aligned with the underlying data structure, thereby enhancing model training fidelity without straying from the original dataset’s essence. Their reliability suggests a broader applicability across various datasets and experimental conditions, emphasizing their significance in achieving stable and predictable results. Delving deeper into these methods, we discern the following characteristics:
• Diversity and Model Generalization: Permutation and Random Permutation, by rearranging time series segments, expose the model to a wider array of patterns, potentially enhancing its generalization capabilities. This strategy effectively simulates real-world conditions where events may occur in varying sequences but still retain their importance.
• Combatting Overfitting: These methods mitigate overfitting by preventing the model from memorizing specific sequences, instead encouraging the recognition of more general patterns.
• Preserving Inherent Data Characteristics: While they alter the sequence, Permutation and Random Permutation maintain the data’s essential features, ensuring the model’s learning remains grounded in the original data’s attributes.
• Primacy of Shape Over Alignment: The RGWs method, utilizing shapeDTW, emphasizes the importance of pattern shape over precise temporal alignment, reflecting real-world scenarios where the overall pattern shape is often more critical than its exact sequence in time.
• Elastic Variations for Robustness: RGWs prepares the model for real-world imperfections by introducing elastic deformations, enhancing robustness to common minor distortions in time-series data.
• Incorporating Domain Knowledge: The RGWs method integrates established practices from time-series analysis by employing shapeDTW, aligning data augmentation strategies with domain-specific knowledge.
Conversely, Rotation, SPAWNER, and Jitter demonstrate more variable outcomes, including underperformance in certain contexts, suggesting their utility may be conditional. Their performance variability underscores the necessity for careful application and calibration, particularly in scenarios demanding data integrity preservation. We also explore the potential reasons for these techniques’ inconsistent performance:
• Hybrid Augmentation Approach: SPAWNER’s complexity, as a hybrid method, lies in its pattern merging and noise introduction, potentially leading to overly complex or unrepresentative patterns.
• Noise: A Delicate Balance: The introduction of noise requires careful moderation to avoid obscuring key data patterns, necessitating a balanced approach for optimal model training.
• Challenges with Inverting Time Series: Rotation may disrupt the natural sequence of time-series data, presenting challenges in contexts where sequence integrity is crucial.
• Potential Ambiguity: Rotation risks creating ambiguity in pattern recognition, potentially complicating classification tasks during model training.
[…]
VI-C Response to RQ2: Influence of Dataset Characteristics on Augmentation Effectiveness and User Recommendations
The realm of time series classification presents perennial challenges, particularly given the heterogeneous nature of datasets. Our experimental phase has shed light on the nuanced responses of various UCR datasets to different augmentation techniques. This discussion aims to unravel the complex interplay between dataset attributes and the efficacy of augmentation strategies.
Figure 14: Accuracy of DA Techniques across Datasets
VI-C1 Dataset Characteristics Analysis
• Length of Time Series: Datasets with extended time series, such as ScreenType, which encompasses 720 units, afford the luxury of more complex augmentations. Techniques like Time Warping and Magnitude Warping, designed for these datasets, adjust the temporal aspects without sacrificing pattern integrity. Conversely, shorter series like ECG200 might find techniques such as Window Slicing less effective, as critical patterns could be lost. Moreover, in datasets like Yoga with a lengthier series (426 units), maintaining the integrity of patterns that resemble image-like structures during any temporal adjustments is paramount.
• Volume of Training Data: Datasets with limited samples, like CBF and ECG200, pose unique challenges due to their restricted diversity. Here, augmentations like Permutation and Random Permutation prove beneficial by providing a wider pattern spectrum, thereby enhancing model robustness and generalization. For instance, Permutation’s notable performance on CBF highlights its role in diversifying smaller datasets. In contrast, larger datasets, such as FordB with 3636 training instances, can withstand more intensive augmentations without the risk of overfitting.
• Dataset Type and Inherent Noise: Dataset specificity significantly sways augmentation selection. For example, Strawberry, categorized under the Spectro type, has spectral patterns sensitive to noise-infused augmentations, evidenced by the lower performance of Jitter in our experiments. Similarly, Yoga, with its image-like formations, necessitates augmentations that avoid excessive noise or distortion, preserving the essential patterns.
• Number of Classes: The class diversity within a dataset also impacts augmentation choices. In a dataset like ECG5000, with five classes, augmentations must accentuate inter-class differences while consolidating intra-class patterns. Our experiments with DGWs and DGW on ECG5000 underscore their proficiency in enhancing class-specific pattern separability, making them suitable for multi-class scenarios.
In addressing RQ2 regarding the influence of dataset characteristics on augmentation effectiveness and user recommendations: Dataset attributes critically dictate the appropriateness and success of data augmentation strategies. The time series length significantly influences augmentation selection, as seen with ScreenType and ECG200, where Time Warping and Window Slicing’s effectiveness varied markedly. The quantity of training data is another decisive factor, as evidenced by Permutation’s varying impact on CBF and FordB. Additionally, the dataset’s nature and its sensitivity to noise influence augmentation compatibility, as demonstrated with Strawberry and Jitter. Furthermore, the class diversity within datasets like ECG5000 necessitates augmentations that bolster class distinctions.
In essence, the careful selection of data augmentation strategies, considering factors like time series length, training data volume, dataset nature, and class variety, is crucial. While transformation-based methods generally exhibit reliable performance, the success of other techniques often depends on specific dataset attributes.
VI-C2 Augmentation Recommendations Based on Dataset Characteristics
VI-C2a Transformation-Based Methods
• General Recommendation: Transformation-Based methods are advisable for datasets where preserving the inherent data structure is critical. These methods maintain the core attributes of the original data, introducing subtle, controlled alterations.
• Specifics: In motion-centric datasets like GunPointAgeSpan, maintaining sequential dependencies is essential, warranting caution with methods like Rotation. For datasets like Strawberry or Yoga, with potentially higher signal-to-noise ratios, Scaling or Time Warping could be beneficial, adding variability without excessive noise. In instances like ScreenType, with its extended time series, transformations adjusting time alignment, such as Time Warping, could effectively introduce meaningful diversity.
• Practical Recommendation: For practitioners dealing with critical pattern-centric time series data, starting with transformation-based methods is prudent. Subsequent experimentation with transformation intensity can help strike a balance between variability and data integrity.
VI-C2b Pattern-Based Methods
• General Recommendation: Pattern-based methods are invaluable for feature-rich datasets or when new structural elements could enhance the dataset.
• Specifics: The efficacy of Random Permutation and Permutation in limited-data scenarios, as seen with CBF, highlights their role in enhancing dataset diversity by reconfiguring existing patterns. In larger datasets like FordB, creating new patterns through methods like SPAWNER can be effective, provided these patterns do not misguide the learning model. In complex datasets like ECG5000, careful pattern blending is essential to avoid over-complicating the learning landscape.
• Practical Recommendation: Practitioners should assess their dataset’s complexity. For data with an already rich feature set, blending might be more beneficial, while in simpler scenarios, it could introduce unwarranted complexity.
VI-C2c Generative Methods
• General Recommendation: Generative methods hold promise, especially for volume-deficient or diversity-lacking datasets. However, ensuring the alignment of generated samples with the original data’s statistical properties is crucial.
• Specifics: Generative methods, such as GAN, have shown potential in datasets like ECG200, where they can replicate complex ECG patterns, enriching the training set. However, for datasets like GunPointAgeSpan, it is imperative that generated samples accurately reflect real-world motion sequences to maintain model relevance. Generative strategies may introduce novel patterns, necessitating careful scrutiny to avoid bias introduction.
• Practical Recommendation: While generative methods offer substantial dataset expansion, quality control of the generated samples is paramount. Ongoing validation against real-world data is recommended.
In summary, our evaluation highlights the consistent performance of Transformation-Based methods, attributed to their respect for the original data characteristics. Conversely, the success of Pattern-Based and Generative Methods may be more context-sensitive, demanding dataset-specific customization and fine-tuning.
VII Limitations and Threats to Validity
Every empirical investigation, irrespective of the meticulousness in its execution, harbors certain limitations. Acknowledging these constraints not only bolsters the study’s credibility but also paves the way for subsequent, more refined research. In this discourse, we deliberate on the inherent limitations of our study and the strategies employed to minimize their implications.
VII-A Internal Validity Threats
VII-A1 Limitation of Data Sources
The primary resource for our study was the UCR time-series classification archive. Although this repository is comprehensive and esteemed within scholarly circles, it encompasses merely a fraction of all conceivable time-series datasets.
• Mitigative Strategy: To counterbalance this limitation, our selection of datasets from the UCR archive was intentionally diverse, spanning various categories of time-series data, from sensor readings to medical diagnostic data.
VII-A2 Scope of Data Augmentation Techniques
Our evaluation encompassed several data augmentation strategies, yet the expanse of existing and emerging methods is vast.
• Mitigative Strategy: Our research prioritized methods that have gained recognition and widespread application. Furthermore, we adopted an inclusive approach by utilizing open-source implementations where feasible, establishing a unified framework conducive to future expansions and adaptations.
VII-A3 Influence of Hyperparameter Configurations
The efficacy of data augmentation strategies is often contingent on hyperparameter configurations. Our adherence to parameters endorsed in scholarly literature does not preclude the potential for alternative configurations to engender different outcomes.
• Mitigative Strategy: We adopted hyperparameters validated by previous empirical studies, ensuring that our findings are anchored in established methodologies and facilitating comparability with extant literature.
VII-A4 Model Specificity
Our analyses relied heavily on deep Convolutional Neural Networks (CNNs), particularly the ResNet architecture. The implications of utilizing alternative models remain an open question.
• Mitigative Strategy: The selection of ResNet was informed by its documented efficacy in time-series classification tasks. We recognize the plethora of alternative models and advocate for additional studies to ascertain the impact of various data augmentation techniques across diverse neural network architectures.
VII-B External Validity Threats
VII-B1 Generalizability Concerns
The applicability of our conclusions beyond the datasets and methods employed in our study may be subject to scrutiny.
• Mitigative Strategy: Our methodological design, characterized by a diverse dataset selection and a comprehensive array of data augmentation techniques, was geared towards maximizing the generalizability of our findings. However, we advise prudence in extending these conclusions to substantially dissimilar scenarios.
VII-B2 Temporal Context of Tools and Methodologies
The machine learning landscape is in a state of perpetual evolution, with constant refinements in tools, methodologies, and best practices.
• Mitigative Strategy: This study captures the state of the art at the time of writing. While our approach was exhaustive given the current landscape, impending advancements may necessitate new interpretations or adaptations of our findings.
VII-B3 Potential Publication Bias
Our literature review may reflect a propensity towards methodologies that have demonstrated success or novelty, inadvertently marginalizing equally potent but underreported techniques.
• Mitigative Strategy: We endeavored for a holistic literature review, although we concede the potential for publication bias. We thus endorse the exploration of both lesser-known methodologies and gray literature in future scholarly endeavors.
In summation, our study, despite its inherent constraints, was conducted with rigorous adherence to scholarly standards. By candidly addressing potential validity threats, we aspire to furnish a robust platform for subsequent inquiries in the realm of time-series data augmentation.
VIII Conclusions and Prospects for Future Research
VIII-A Concluding Remarks
This investigation embarked on a meticulous examination of data augmentation strategies, specifically curated for time-series classification. Upon an exhaustive retrospection spanning the past 10 years, we discerned a plethora of techniques, amassing over 60 distinct methods from in excess of 100 scholarly articles. These methods were subsequently assimilated into an innovative taxonomy, crafted with a focus on time-series classification. This taxonomy segregates the techniques into five quintessential categories: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation, thereby providing a structured reference point for scholars and industry practitioners.
Our inquiry transcended theoretical postulation, venturing into pragmatic assessments. We applied an array of 15 prevalent data augmentation strategies to 8 categories of datasets derived from the esteemed UCR time-series classification archive. The decision to employ ResNet as the evaluative model was strategic, motivated by its documented success in analogous research contexts, thereby ensuring a solid experimental groundwork. Our multi-faceted evaluation approach, encompassing Accuracy, Method Ranking, and Residual Analysis, yielded insights into the nuanced performance dynamics of each technique.
The experimental phase spanned approximately 2 weeks, culminating in a baseline model accuracy of 88.94 ± 11.83%. Notably, the Permutation method (boasting an average rank of 4.125 and an accuracy of 89.20 ± 11.95%, and prevailing in 2 datasets) enhanced the model’s performance substantially, whereas the Rotation method (84.75 ± 12.92%) was detrimental. Other strategies, including Window Warping and GAN, manifested negligible deviations, underscoring the situational specificity of these methods. Transformation-Based Methods, in particular, have consistently exhibited their merit, potentially attributable to their finesse in maintaining and subtly modifying the inherent structures of time-series data.
Our analysis, extended across various UCR datasets, revealed that the characteristics intrinsic to each dataset significantly influence the success of augmentation strategies. For instance, datasets with extended durations, such as ScreenType, are amenable to temporal transformations, whereas datasets with multiple classifications, like ECG5000, necessitate distinct augmentations for each class. Our recommendations advocate for Transformation-Based strategies when data integrity is paramount, and Pattern-Based strategies for datasets abundant in features. Generative strategies, albeit potent, demand meticulous alignment with the original data characteristics. Predominantly, Transformation-Based strategies exhibit a superior adaptability across diverse contexts.
To contribute tangibly to the field, we have integrated these strategies into a comprehensive Python library. This utility, engineered for user-friendliness, simplifies the application of diverse augmentation strategies, thereby catalyzing further experimental endeavors and innovations.
In essence, this research presents an integrative perspective on the contemporary landscape of data augmentation for time-series classification, melding theoretical frameworks with empirical evidence. The revelations and resources introduced herein are positioned to catalyze continued progress in this domain, fortifying machine learning models against the challenges posed by data limitations, and enhancing their generalizability and robustness.
VIII-B Future Research Trajectories
The domain of data augmentation for time-series classification is expansive and continually evolving. The milestones achieved in this research constitute a foundational layer, with expansive potential for further exploration and refinement. Outlined below are prospective directions for subsequent inquiries.
VIII-B1 Exploration of Diverse Neural Network Architectures
While the present study was anchored around a select group of established neural network models, the dynamic nature of deep learning architectures intimates the potential advantages of a more extensive exploration. Particularly, the adaptability of emerging models, especially those designed for sequential data like Transformer models, warrants comprehensive investigation.
VIII-B2 Extension to an Expanded Dataset Spectrum
The scope of our study was limited to a specific selection of datasets, representing only a segment of the myriad applications of time-series classification. Future research could broaden this scope to encompass datasets from diverse fields such as finance, environmental science, or social dynamics, thereby testing the universality and domain-specific nuances of the augmentation techniques.
VIII-B3 Advanced Augmentation Selection and Synthesis
The strategic amalgamation of various augmentation techniques holds the promise of enhancing their individual efficacies. However, indiscriminate combinations may induce data redundancy or inefficiencies. Innovative methods like RandECG [99], which judiciously integrate multiple augmentations through transparent, interpretable hyperparameters, herald a promising future for augmentation synthesis. Adapting such strategies for time-series data, potentially with the incorporation of advanced frameworks like reinforcement learning or meta-learning, could optimize augmentation processes. Furthermore, ensuring the computational efficiency of these algorithms, especially in scenarios with extensive datasets or limited processing capabilities, remains an essential consideration.