### v5

# Data Augmentation for Time-Series Classification: a Comprehensive Survey

Abstract
Data Augmentation (DA) has become a critical approach in Time Series Classification (TSC), primarily for its capacity to expand training datasets, enhance model robustness, introduce diversity, and reduce overfitting. However, the current landscape of DA in TSC is plagued with fragmented literature reviews, nebulous methodological taxonomies, inadequate evaluative measures, and a dearth of accessible and user-oriented tools. This study addresses these challenges through a comprehensive examination of DA methodologies within the TSC domain. Our research began with an extensive literature review spanning a decade, revealing significant gaps in existing surveys and necessitating a detailed analysis of over 100 scholarly articles to identify more than 60 distinct DA techniques. This rigorous review led to the development of a novel taxonomy tailored to the specific needs of DA in TSC, categorizing techniques into five primary categories: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation. This taxonomy is intended to guide researchers in selecting appropriate methods with greater clarity. In response to the lack of comprehensive evaluations of foundational DA techniques, we conducted a thorough empirical study, testing nearly 20 DA strategies across 15 diverse datasets representing all types within the UCR time-series repository. Using ResNet and LSTM architectures, we employed a multifaceted evaluation approach, including metrics such as Accuracy, Method Ranking, and Residual Analysis, resulting in a benchmark accuracy of 84.98 ± 16.41% in ResNet and 82.41 ± 18.71% in LSTM. Our investigation underscored the inconsistent efficacies of DA techniques, for instance, methods like RGWs and Random Permutation significantly improved model performance, whereas others, like EMD, were less effective. Furthermore, we found that the intrinsic characteristics of datasets significantly influence the success of DA methods, leading to targeted recommendations based on empirical evidence to help practitioners select the most suitable DA techniques for specific datasets. To improve practical use, we have consolidated most of these methods into a unified Python Library, whose user-friendly interface facilitates experimenting with various augmentation techniques, offering practitioners and researchers a more convenient tool for innovation than currently available options.
Index Terms:
Data Augmentation, Time-Series Classification, Machine Learning, Deep Learning.
I Introduction
In machine learning, the pursuit of robust model performance often depends on the availability of extensive datasets, a challenge exacerbated by the frequent scarcity of data [1]. Data augmentation addresses this issue by generating additional, analogous data to expand limited datasets, thereby mitigating data insufficiency constraints. This technique, renowned for its cost-effectiveness and efficacy, has gained immense traction in various machine learning disciplines, significantly improving model accuracy in environments characterized by limited data. For instance, in computer vision, image augmentation is relatively straightforward, with a variety of simple yet effective techniques capable of producing discernible images even when subjected to noise or partial cropping [2].
In contrast, time-series data augmentation has received less attention. Existing strategies often overlook the intrinsic properties defining time-series data, such as temporal correlation, a critical aspect that distinguishes it from other data types. The potential for transforming time-series data into frequency or time-frequency domains further complicates the application of conventional augmentation techniques, particularly those borrowed from image or speech processing, leading to potential inaccuracies in synthetic data generation. Additionally, the effectiveness of data augmentation methods is highly task-dependent, requiring a nuanced approach that addresses the unique demands of tasks like forecasting and classification within the time-series context.
This paper offers an extensive examination of the current state of Data Augmentation (DA) for Time Series Classification (TSC), exploring a wide range of techniques and methodologies that have shaped the field over the past decade, providing readers with a comprehensive understanding of its evolution.
- Comprehensive Survey and Novel Taxonomy: To bring clarity and structure to the diverse array of techniques in this field, we introduce a novel taxonomy specifically designed for DA in TSC. This classification system organizes methods into five fundamental categories: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation, enabling a systematic and streamlined approach to selecting and analyzing DA methods for TSC.
- Detailed Evaluation: Beyond a traditional survey and taxonomy, this study makes a significant contribution by rigorously evaluating nearly twenty DA techniques for TSC. These methods were tested using two established neural network models, providing a robust performance benchmark. Our empirical analysis offers a comprehensive perspective on the relative strengths and weaknesses of these techniques, delivering valuable insights for both academia and industry.
- Unified Library: To complement our efforts, we have developed a Python Library that encapsulates most of the aforementioned time-series data augmentation algorithms, designed for ease of implementation. This initiative addresses the shortcomings of existing tools, offering a more user-friendly alternative.
The insights gained from our research provide a broad view of current techniques while highlighting potential areas for improvement and future research. We aim to foster further innovation in DA for TSC, paving the way for the development of more sophisticated and effective methods tailored to the complexities of time-series data. The main contributions of this manuscript are as follows:
  • We conduct a thorough review of advancements in Data Augmentation for Time Series Classification (DA for TSC) over the past decade;
  • We advocate a novel, purpose-built taxonomy for DA for TSC, segmenting techniques into quintessential categories: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation;
  • We perform an in-depth evaluation of nearly twenty DA for TSC techniques, utilizing two renowned neural network models, and offer an exhaustive discourse on the findings;
  • We introduce a Python Library of preeminent time series data augmentation algorithms, offering a unified, user-friendly platform for the application of diverse data augmentation strategies to time series datasets.
The structure of this article is meticulously organized to facilitate a comprehensive understanding of our investigation, as depicted in Figure 1. Section II embarks on a thorough literature review, delineating the state of the art in Data Augmentation (DA) for Time Series Classification (TSC), establishing the foundation upon which our study is predicated. In Section III, we elaborate on the methodological framework adopted for our survey, including the criteria for selection and analysis of relevant works, and engage in a detailed discussion of the findings derived therefrom. Section IV introduces a novel taxonomy for DA techniques, methodically categorizing these methodologies into five distinct principal categories, thereby offering a structured lens through which the current landscape can be viewed. Section V is dedicated to the elucidation of our experimental design, meticulously detailing the setup employed in our empirical study and articulating the research questions poised at the outset. Subsequently, Section VI addresses these research questions, presenting a critical analysis of the experimental outcomes and drawing insights pertinent to the field of DA for TSC. Section VII ventures into an exploration of the limitations inherent to our study and the potential threats to the validity of our findings, fostering a transparent discourse on the scope and applicability of our conclusions. The article culminates in Section VIII, where we synthesize our contributions to the domain and outline directions for future research.
Figure 1: The Overall Organization of this Survey Paper
II Background and Fundamentals
Time Series Classification: The domain of Time Series Classification (TSC) represents a pivotal frontier in data mining, dedicated to the categorization of time series into distinct classes predicated upon historical pattern analysis [3]. Inherent in a time series is a sequence of data points, chronologically ordered and indexed over regular intervals, underscoring the temporal dynamics integral to various practical realms, including health diagnostics, financial market analytics, among others. The intricate nature of time series data, marked by heterogeneous lengths, noise interferences, and nonlinear trajectories, poses substantial analytical challenges.
Data Augmentation: Within the spheres of machine learning and deep learning, data augmentation stands out as a strategy for amplifying the training dataset by generating modified iterations of the original data instances [4]. This process, applicable across diverse tasks such as image recognition, involves transformations such as rotation, scaling, or translation. In the context of TSC, augmentation encompasses techniques like noise injection, slicing, jittering, and scaling, serving dual core objectives: enhancing model generalization to diminish overfitting and addressing class imbalances within datasets.
Despite its established efficacy in various machine learning applications, data augmentation’s integration into TSC remains a vibrant research avenue, necessitating deeper exploration. Historically, data augmentation has proven indispensable in scenarios plagued by data scarcity or imbalance, as evidenced in applications like AlexNet for ImageNet classification [5]. Recent scholarly endeavors have extended these techniques to time series data, albeit with modifications to accommodate the data’s unique characteristics [6, 7, 8, 9].
The singular nature of time series data often precludes the direct application of standard augmentation methods, necessitating adaptations or entirely novel methodologies, particularly for tasks like classification, forecasting, and anomaly detection within time series [10]. The past decade has witnessed a proliferation of TSC algorithms, spurred by the increased availability of time-series data and its ubiquity across tasks necessitating human cognitive processes [11, 12]. The diverse dataset types within the UCR/UEA archive further testify to the expansive applicability of TSC [13].
Addressing data imbalances, a prevalent issue in machine learning tasks, requires strategic data augmentation interventions, such as oversampling or undersampling, to rectify biases towards majority classes and enhance minority class performance. Techniques like SMOTE (Synthetic Minority Over-sampling Technique) generate synthetic examples to balance class distributions, while their counterparts reduce majority class instances [14]. However, these approaches necessitate careful calibration to avoid overfitting or information loss.
Innovative contributions to the field include Guennec et al.’s two-layer convolutional network for TSC [15], Germain et al.’s time series-specific augmentation methods [16], and Pfister et al.’s exploration of wearable sensor data augmentation [6]. Haradal et al. ventured into biosignal classification, employing generative adversarial networks (GAN) for synthetic time series data generation [17]. Fons et al. examined financial time series classification, evaluating various data augmentation strategies with deep learning models [18]. Nonaka et al. proposed ECG-specific data augmentation, enhancing atrial fibrillation classification using single-lead ECG data without modifying the DNN architecture [19]. Goubeaud et al. introduced White Noise Windows, a novel time series data augmentation technique, in 2021 [20].
Despite its advantages, data augmentation is not without its challenges, including the risk of overfitting and the potential ineffectiveness with high-dimensional or complex data structures [4, 21].
II-A Relevant Survey & Study
The scholarly landscape is replete with surveys on data augmentation, spanning domains from image processing [22] to natural language processing [23], though comprehensive studies in the time series sector remain scarce. Iwana et al. conducted an empirical assessment of time series data augmentation methods using neural networks, albeit with a focus limited to classification and excluding frequency domain or externally trained methods [4]. Wen et al. expanded upon this work, incorporating time series forecasting and anomaly detection, though their study utilized a relatively constrained dataset from the Alibaba Cloud monitoring system [21].
Although these studies offer valuable insights into data augmentation techniques, there is a conspicuous absence of a holistic view that encompasses recent advances and diverse application fields. Several data augmentation techniques referenced in the literature are either antiquated or no longer align with contemporary research and practice standards, thereby failing to offer a user-friendly and accessible toolset for researchers and practitioners. This article endeavors to address this gap in the scholarly discourse by conducting a comprehensive analysis of the current landscape of data augmentation within the domain of time-series classification. It aims to serve as a pivotal reference point, guiding future research directions and fostering innovation in this rapidly evolving field.
III Survey Methodology
This paper adopts a tripartite research methodology encompassing Survey, Evaluation, and Experimental Study Setup. The Survey delineates our systematic approach to literature accumulation and analysis. The Evaluation introduces a novel taxonomy of time series data augmentation, predicated on distinctive characteristics. The Experimental Study Setup discusses the authoritative datasets employed, evaluation metrics, and potential variables influencing outcomes, thereby ensuring a rigorous analysis of time series data augmentation methodologies.
III-A Paper Collection Criteria of the Survey
The literature aggregation for this study adheres to stringent inclusion criteria, encompassing:
  • Papers elucidating the fundamental concepts of data augmentation for time-series classification or its associated facets.
  • Studies proposing enhancements to existing data augmentation methodologies, tools, or frameworks within the realm of time-series classification.
  • Contributions introducing benchmarks or datasets tailored for data augmentation in time-series classification.
  • Literature offering comprehensive reviews or surveys on the contemporary landscape of data augmentation for time-series classification.
Exclusions pertain to studies applying relevant techniques outside the time-series context and those focusing on time-series classification, forecasting, or anomaly detection without integrating data augmentation methodologies.
III-B Paper Collection Methodology of the Survey
Our literature retrieval process commenced with exhaustive searches across renowned scientific repositories like Google Scholar and arXiv, utilizing precise keyword combinations. Recognizing the interdisciplinary application of time series data augmentation, we adopted a snowballing strategy for comprehensive coverage [24]. The accumulated literature was subsequently categorized as Survey/Review papers, Research papers, or Methods papers, following established academic conventions [25, 26, 27].
The keywords used for searching and collecting are listed below. [DA for TSC] means Data Augmentation for Time-Series Classification, which is the focus of this research project. The search strategy encompassed several iterations, as detailed below:
  • DA for TSC + Survey / Review
  • DA for TSC + Techniques / Methods
  • DA for TSC + Evaluations / Comparison
III-C Result and Discussion of the Survey
TABLE I: Paper Query Results
Paper Query Results Key Words Hits Title
[DA for TSC] + Survey / Review 25 20 8
[DA for TSC] + Techniques / Methods 273 102 72
[DA for TSC] + Evaluations / Comparison 90 55 26
Query - - 106
Snowball - - 41
Overall - - 147
The terms “Key Words”, “Hits” and “Title” refer to the number of papers returned by the search, the number of papers that remain after title framing and the number of papers that remain after body framing, respectively. The search strategy yielded 147 relevant publications until 10th June, 2024, with a significant proportion originating from keyword searches and the remainder from snowballing. Trends in annual publication indicate a growing interest in this research area, particularly after 2019, as shown in Figure 2.
Figure 2: Annual Publication Trends in Data Augmentation for Time-Series Classification
The distribution of publications across various publishers underscores the dominance of IEEE and arXiv, with significant contributions also emanating from Elsevier, MDPI, ACM, and Springer.
Figure 3: Publisher Distribution of Publications in Data Augmentation for Time-Series Classification
A geographical analysis of the publications reveals a preponderance of contributions from the United States, followed by China, Japan, Germany, and the United Kingdom, with disparate representation from other global regions.
Figure 4: Publication Count by Country in Data Augmentation for Time-Series Classification
An industry-centric examination, based on the North American Industry Classification System (NAICS), highlights a predominant application within the Information Technology sector, followed by Healthcare, Agriculture, Astronomy, Finance, and Construction, indicating the methodologies’ broad applicability.
Figure 5: Publication Count by Industry in Data Augmentation for Time-Series Classification
IV Data Augmentation for Time-Series Classification Techniques
The taxonomy presented in this study for data augmentation techniques applicable to time-series classification (TSC) emerges from an exhaustive literature survey and rigorous research. It compartmentalizes the techniques into five principal categories: Transformation-Based Methods, Pattern-Based Methods, Generative Methods, Decomposition-Based Methods, and Automated Data Augmentation Methods. This categorization, meticulously refined and expanded upon from preceding studies [10, 4, 21], lays the groundwork for our in-depth exploration and experimentation.
Several factors underpin the formulation of this taxonomy. Primarily, it encompasses a broad spectrum of contemporary methods, thereby paving the way for domain-centric experimental undertakings in specialized segments like the time and frequency domains. Moreover, it offers an organized synopsis of the myriad techniques used for augmenting time-series data, thereby simplifying the comparative analysis and comprehension of diverse methods. Such a structured approach is indispensable for researchers and practitioners alike, guiding them in the judicious selection of data augmentation techniques tailored to their specific requirements. Furthermore, by highlighting the existing lacunae and prospects in the current scenario, it spurs the innovation of novel techniques.
The distinct categories within this taxonomy epitomize different methodologies in data augmentation. Transformation-Based Methods involve the alteration of original data via various transformations. Pattern-Based Methods, on the other hand, generate new instances through the extraction and subsequent recombination of patterns inherent in the original data. Generative Methods utilize generative models to synthesize new data instances. Decomposition-Based Methods, in contrast, entail the disintegration of the original time-series into multiple components, which are then individually augmented and reassembled. Automated Data Augmentation Methods streamline the augmentation process, often leveraging data-driven insights to ascertain the most efficacious augmentation strategy.
The ensuing subsections provide an exhaustive discourse on each category, elucidating their respective methodologies, merits, and potential use-cases. This detailed analysis serves dual purposes: guiding users in the selection of appropriate data augmentation strategies and fostering the conception of avant-garde methods in forthcoming research endeavors.
{forest}
[…]
for Time-Series
Classification, align=center, before computing xy=l=2cm [Transformation-Based
Methods, align=center [Magnitude [Jittering [6, 28, 29, 30, 31, 32, 33]] [Rotation [6, 34, 32, 33]] [Scaling [6, 7, 35, 29, 30, 32]] [Magnitude Warping [6, 35, 36, 30]] [White Noise Windows [20]] ] [Time [Permutation [6, 34, 30]] [Slicing [15, 29]] [Time Warping [35, 32]] [Window Warping [15, 37, 7]] [Window Slicing [15]] ] [Frequency [Frequency Warping [VTLP [38]] [SFCC [39]] ] [Frequency Slicing [29]] [Random Noise Boxes [40]] ] [Time-Frequency [Frequency Transform [STFT [41]] ] [Surrogate Data Augmentation [42]] ] ] [Pattern-Based
Methods, align=center [Magnitude [TS-DENSE [43]] [RIM [44]] [SPAWNER [45, 46]] [ECGAug [47]] ] [Time [Guided Warping [48] [RGW] [RGWs] [DGW] [DGWs] ] [Interpolation [Interpolation [49]] [Shapelets-based SMOTE [14]] [DTW-SMOTE [50]] ] [DBA [16, 51]] [ASP [52]] [Mapping Label [53]] [StyleTime [54]] [DTW-MERGE [55]] ] [Frequency [EMDA [56]] [SFM [57]] ] [Time-Frequency [Sample Combining [58]] [Audio Augmentation [59, 60]] ] ] [Generative
Methods, align=center [Statistical Models [Gaussian trees [61]] [GRATIS [62]] ] [Deep Learning, l=2cm [GAN [GAN [17, 63, 64, 36, 65, 66]] [Wasserstein GAN [67]] [Conditional GAN [68, 69]] [EmotionalGAN [70]] [BiLstM-CNN GAN [71]] [Time Series GAN [72, 73, 74, 75, 76]] [GRU-CNN GAN [77]] [Imputation Balanced GAN [78]] [PlethAugment [79]] ] [Encoder Decoder [AE [80, 81]] [VAE [82, 83, 84, 85, 86, 87]] ] [ODENet [88]] ] [Dynamical System [89, 90]] [Signal Augmentation [91]] [IMDA [92]] ] [Decomposition-Based
Methods, align=center [EMD [93, 94]] [STL [95, 96, 97, 98]] ] [Automated Data
Augmentation
Methods, align=center [RandECG [99]] [Fast AutoAugment [82]] [Multiple Approaches [100]] [Hybrid Framework [101]] [JobDA [102]] [Adaptive Weighting Scheme [103]] [MODALS [104]] ] ]
Figure 6: Taxonomy of Techniques of Data Augmentation for Time-Series Classification
IV-A Transformation-Based Methods
Transformation-based methods hinge on the principle of applying an array of transformations to the original time-series data. The objective is to retain the fundamental structure while introducing nuanced variations, thereby bolstering model resilience. These methods, characterized by their simplicity and intuitiveness, typically encompass elementary mathematical operations, making them versatile across diverse data types and domains. The scope of transformations extends to various data facets, culminating in four distinct sub-categories.
Figure 7: Illustrative Examples of Transformation-Based Methods on the ECG5000 Dataset (Generated instances are depicted in orange)
IV-A1 Magnitude Transformations
Magnitude transformations alter the amplitude of time-series data, an approach especially pertinent in contexts where the data’s magnitude is informative, such as signal processing or audio analysis.
For instance, the study by Terry et al. [6] delves into data augmentation techniques like Jittering, Scaling, Rotation, Magnitude-Warping, etc, applied for Parkinson’s disease movement classification using wearable sensor data. Research underscores that an amalgamation of DA methods markedly propels classification efficacy, jumping from a baseline of 77.52% to 86.88%. This finding signals a promising trajectory for wearable sensor data augmentation. Similarly, Shuang et al. [35] harnesses Scaling, Magnitude Warping, and other techniques on I/Q signal data, culminating in a peak recognition accuracy of 93.68% and successful differentiation between QAM16 and QAM64 signals. Additionally, Goubeaud et al. [20] introduce White Noise Windows (WNW), a novel augmentation technique that fortifies classification methods and mitigates overfitting.
IV-A2 Time Transformations
Time transformations adjust the temporal elements of the data, such as data point sequencing or timing. They are instrumental when the data’s temporal dynamics are crucial, as seen in tasks like time-series classification, forecasting, or sequence prediction.
In addition to the aforementioned magnitude transformations, the work of Terry et al. [6] also integrates time-based data augmentation techniques such as Permutation and Time-Warping, further demonstrating their utility in enhancing Parkinson’s disease movement classification through wearable sensor data. Similarly, the study by Shuang et al. [35] employs Time-Warping among its suite of techniques, showcasing its effectiveness in improving the recognition accuracy of I/Q signal data. These instances underline the broad applicability and potential of time transformations in augmenting data for varied classification tasks.
In the realm of construction equipment IMU data, Rashid et al. [7] innovates with Window-Warping (WW), a technique that synthesizes training data, thereby enhancing machine learning model proficiency in activity recognition tasks, with K-Nearest Neighbors (KNN) witnessing substantial improvements. Furthermore, Guennec et al. [15] discusses two augmentation methods for time series classification with Convolutional Neural Networks: Window Slicing (WS) and Window Warping (WW). These strategies, coupled with a semi-supervised Dataset Mixing technique, bolster classification performance, particularly in smaller datasets.
IV-A3 Frequency Transformations
Frequency transformations tinker with the time-series data’s frequency components, a strategy quintessential in scenarios where the data’s frequency content is vital, such as in spectral or vibration analysis.
In 2023, Yang et al. [39] pioneers Stratified Fourier Coefficients Combination (SFCC), a method that merges original Fourier coefficients to augment time series datasets, thereby boosting deep neural network performance. This efficient technique achieves state-of-the-art results. In a similar vein, Goubeaud et al. [40] unveils Random Noise Boxes (RNB), an innovative augmentation technique optimized for spectrogram classification. While particularly potent for ECG and sensor datasets, its efficacy diminishes for motion datasets, underscoring the need for vigilant application to avert performance deterioration.
IV-A4 Time-Frequency Transformations
Time-frequency transformations concurrently tweak both temporal and frequency aspects. This dual approach is advantageous when both elements are pivotal, enabling the generation of new instances that preserve the original data’s temporal and frequency structures.
In 2019, Lee et al. [42] employs data augmentation techniques, notably Amplitude Adjusted Fourier Transform (AAFT) and Iterative Amplitude Adjusted Fourier Transform (IAAFT), to concoct surrogate data conducive to deep learning. These strategies maintain the time-domain statistics and power spectrum of the original time series, thereby enriching the training of pre-trained Convolutional Neural Networks (CNNs). The research corroborates the efficacy of these methods in automated rehabilitation classification tasks, suggesting avenues for further exploration.
IV-B Pattern-Based Methods
Pattern-based methods concentrate on discerning and capitalizing on specific patterns or structures inherent in the time-series data. These patterns, whether local or global, simple or complex, often involve machine learning models for identification and utilization. The hallmark of this category is its emphasis on the data’s intrinsic patterns, a focus that proves invaluable in contexts with pronounced structural or pattern-based components, such as electrocardiograms (ECG) or stock market data.
Figure 8: Illustrative Examples of Pattern-Based Methods on the Strawberry Dataset (Generated instances are depicted in orange)
IV-B1 Magnitude Pattern Mining
This niche capitalizes on significant local structures or motifs within the data, exploiting these elements to spawn new, realistic data instances.
In 2019, Kamycki et al. [45] introduces SPAWNER, an innovative data augmentation method that aligns input sequences and restricts the warping path to produce synthetic examples, thereby enhancing the performance of time-series classifiers on benchmark datasets. In a parallel study, Aboussalah et al. [44] employs Recurrent In-fill Masking (RIM) in 2022, a novel augmentation method demonstrated to bolster time series classification task performance across synthetic and real-world datasets, outperforming traditional training and other augmentation strategies.
IV-B2 Time Pattern Mining
These methods adjust the data’s temporal characteristics, maintaining the original data’s temporal structure while generating new instances.
Forestier et al. [16] advocates for a data augmentation method that utilizes Dynamic Time Warping (DTW) to synthesize examples by averaging time series sets, a method especially potent in scenarios with limited example availability. The study confirms that augmenting full datasets typically enhances accuracy, with notable improvements in 56 of the 85 tested datasets. Similarly, Iwana et al. [48] and Cheolhwan et al. [49] introduce novel methods that leverage time warping and cubic interpolation, respectively, to generate additional training samples effectively, thereby optimizing model performance across various datasets.
IV-B3 Frequency Pattern Mining
Frequency Pattern Mining methods focus on manipulating the frequency components of time-series data. These methods are pivotal when the data’s frequency content is crucial, enabling the generation of new instances that preserve the original data’s frequency structure.
In [56], Loris et al. introduce a technique known as EMDA, or Equalized Mixture Data Augmentation, a method premised on spectrograms. EMDA operates by forging a new spectrogram through the weighted average of two randomly selected spectrograms from the same class, incorporating perturbations like time delay and equalizer functions. This strategy, as per the authors, enhances the diversity and robustness of audio signals, thereby fortifying classification performance. Conversely, Cui et al. [57] advocates for Stochastic Feature Mapping (SFM), a data augmentation method tailored for deep neural network acoustic modeling. SFM, a label-preserving transformation, statistically transmutes one speaker’s speech data into another’s within a specified feature space. The essence of SFM is to amplify pattern variations via transformations, thereby bolstering the neural networks’ classification invariance and generalization capabilities.
IV-B4 Time-Frequency Pattern Mining
Time-Frequency Pattern Mining methods adjust both time and frequency components of the data. This dual-pronged approach is beneficial when both the data’s temporal dynamics and frequency content are integral, as seen in domains like speech recognition or music analysis.
In 2020, Zhao et al. [58] proposes a unique method that transitions time-domain data into the frequency domain via a discrete cosine transform (DCT), subsequently generating artificial data by merging different frequency bands. These synthesized signals are then reverted to the time domain through an inverse DCT, with experimental outcomes underscoring the method’s efficacy. In a related study, Nanni et al. [59] employs data augmentation strategies to enhance the classification accuracy of animal sounds using convolutional neural networks (CNNs). These strategies, encompassing transformations of audio signals and their spectrograms, utilize techniques like pitch shift, noise addition, and equalized mixture. The findings affirm that data augmentation bolsters the accuracy and generalization capabilities of CNNs, with a fusion of different CNNs yielding optimal results.
IV-C Generative Methods
Generative Methods constitute a separate class within the taxonomy of time-series data augmentation techniques. These methods leverage generative models to craft new time-series instances, diverging from transformation-based and pattern-based methods by incorporating a learning phase. During this phase, the model discerns the data’s underlying distribution, subsequently spawning new instances from this distribution. This category is invaluable when the objective is to generate diverse, realistic synthetic instances that mirror the original data’s statistical attributes.
Figure 9: Illustrative Examples of Generative Methods (GAN) on the FordB Dataset (Generated instances are depicted in orange)
IV-C1 Statistical Models
Statistical Models utilize generative models with pre-established structures to produce new time-series instances. These models are advantageous when the data’s underlying distribution is known or can be reasonably inferred, facilitating the generation of new instances that adhere to the original data’s distribution.
In [61], a Parsimonious Mixture of Gaussian Trees (2MoGT) model is presented by Cao et al., designed for data augmentation in scenarios with imbalanced and multimodal time-series classification. This model employs Chow-Liu trees to encapsulate dependencies among time-series features and a Gaussian mixture model to accommodate multimodality. It surpasses alternative oversampling methods, especially in datasets with multimodal minority class distributions, and minimizes storage complexity. Additionally, Kang et al. [62] introduces GRATIS, a data augmentation technique that generates time series with adjustable characteristics. Utilizing a mixture autoregressive (MAR) model, attuned to specific features, it spawns diverse time series. The resulting series exhibit diversity, attributable to the mixtures’ adaptability and random parameter configurations, proving especially beneficial for analyzing collections of time series with varying lengths, scales, and characteristics.
IV-C2 Deep Learning Models
Deep Learning Models employ generative models without predefined forms to create new time-series instances. These models are beneficial when dealing with data containing complex, non-linear, or multi-dimensional patterns, as they can generate new instances that encapsulate these intricate patterns.
In [17], a data augmentation strategy based on Generative Adversarial Networks (GANs) is presented by Haradal et al. for biosignal classification. This method incorporates a Recurrent Neural Network (RNN) with LSTM units for the generation of time-series data. Its efficacy is validated by an uptick in classification accuracy, underscoring its viability for synthesizing biosignals and enhancing biosignal classification. Conversely, in 2022, Garcia et al. [64] applies GANs for data augmentation, synthesizing light curves from variable stars. This methodology, merged with an innovative resampling technique and evaluation metric, markedly elevates the classification accuracy of variable stars, proving particularly potent for unbalanced datasets and in pinpointing GAN-overfitting scenarios.
Several researchers have also harnessed specialized GANs for data augmentation. For instance, Ehrhart et al. [69] employed a conditional Generative Adversarial Network (cGAN) with a diversity term for data augmentation, utilizing an LSTM for the generator and an FCN for the discriminator. This method synthesized moments of stress from physiological sensor data, bolstering classifier performance and emerging as a promising technique for stress detection in wearable physiological sensor data. Moreover, in 2023, Yang et al. [72] introduced a time-series Generative Adversarial Network (TS-GAN) for augmenting sensor-based health data, generating synthetic time-series data that preserves the original data’s temporal dependencies, thereby enhancing model performance. The study asserts that TS-GAN surpasses conventional augmentation methods in maintaining data consistency. Additionally, Deng et al. [78] implemented a novel data augmentation strategy, IB-GAN, merging imputation and balancing techniques. It boosts classification performance by synthesizing synthetic samples for under-represented classes, averting the mode collapse often seen in standard GANs. This approach records substantial performance enhancements against cutting-edge baselines. In [70], an emotion state identification framework and ECG signal augmentation algorithm based on GAN are proposed by Chen et al. EmotionalGAN is utilized to synthesize synthetic ECG data, which is then amalgamated with the original training set, enhancing emotion state classification. The findings indicate a notable improvement in classification performance with the integration of EmotionalGAN-synthesized data.
Models based on the Encoder Decoder structure have also gained traction in Data Augmentation for Time-Series Classification. Moreno et al. [85] employs Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) for data augmentation, enhancing prediction accuracy on small datasets, with substantial gains realized through computational effort. The study also introduces alterations to balance class representation in synthetic samples. Zha et al. [80] unveils a straightforward yet potent self-supervised model for time series generation, termed the Masked Autoencoder with Extrapolator (ExtraMAE). The synthetic time series crafted by ExtraMAE exhibit remarkable practicality and fidelity, eclipsing all benchmarks while remaining more compact and expedient in a minimalist design. Furthermore, Hsu et al. [87] introduces two Variational Autoencoder (VAE)-based data augmentation strategies for unsupervised domain adaptation in Automatic Speech Recognition (ASR). These methods, encompassing nuisance attribute replacement and soft latent nuisance subspace perturbation, alter nuisance attributes of speech, achieving approximately a 35% absolute Word Error Rate (WER) reduction.
IV-D Decomposition-Based Methods
Decomposition-Based Methods disassemble the original time-series into multiple components, augment each component independently, and subsequently reassemble them to produce new instances. This unique tactic of deconstructing the time-series data and individually addressing each component proves advantageous when various components of the data necessitate distinct augmentation techniques.
Figure 10: Illustrative Examples of Decomposition-Based Methods (EMD) on the FordB Dataset (Generated instances are depicted in orange)
In [93], a data augmentation strategy predicated on Empirical Mode Decomposition (EMD) and an Integrated LSTM neural network is showcased by Li et al. Raw hand motor signals are fragmented by EMD, followed by the augmentation of the decomposed signals through LSTM. The method, trialed on the Ninaweb public dataset, demonstrated a 5.2% ascension in classification accuracy compared to benchmark outcomes. In 2019, Wen et al. [96] introduces RobustSTL, a pioneering algorithm for time-series decomposition that tackles anomalies, sudden trend shifts, seasonality alterations, and noise. It employs bilateral filtering for noise eradication, LAD regression for trend extraction, and non-local seasonal filtering for seasonality determination. The algorithm surpasses existing solutions in managing complex time-series data. An extension of RobustSTL is presented in [98] by Wen et al., designed to accommodate multiple seasonalities in time series data, utilizing a specialized generalized ADMM algorithm. This method outperforms leading algorithms like MSTL, STR, TBATS in terms of effectiveness and computational efficiency, with future explorations geared towards multi-resolution seasonal-trend decomposition for further enhancements.
IV-E Automated Data Augmentation Methods
Automated Data Augmentation Methods strive to automate data augmentation processes, often discerning the optimal augmentation strategy directly from the data. This category distinguishes itself with its focus on automation and adaptability, ideal for circumstances where manual selection and adjustment of augmentation techniques are either impractical or not optimal.
In 2022, Park et al. [82] harnesses PoseVAE and Fast AutoAugment for data augmentation, with PoseVAE rejuvenating skeleton data based on initial human poses, thereby generating stable and novel data. Fast AutoAugment fine-tunes the augmentation policy employing RNN-based reinforcement learning. The research concludes that these methodologies can effectively augment skeleton data, boosting model performance.
[…]
It fine-tunes four universal data transformation operations to conform to various data modalities, creating challenging examples and employing additional loss terms to refine label-preserving transformations. It manifests effectiveness across diverse modalities, including text, tabular, time-series, and images.
V Experimental Study Setup
In this section, we detail the comprehensive comparative evaluations performed, using various data enhancement techniques. These evaluations are designed to provide an empirical comparison between different data augmentation strategies. The primary focus of the comparison is to assess the influence of these methods on the performance of neural network models, especially in the context of time series datasets. This approach allows for a deeper understanding of how these techniques can be optimized to enhance model performance, offering a broader view of their potential impact across various applications.
V-A Datasets
This research encompasses the application of several time series classification techniques to various datasets from the University of California, Riverside (UCR) time series classification repository [13]. The UCR repository is a rich collection of time series datasets, extensively utilized within the academic community for benchmarking time series classification algorithms. For this study, fifteen representative categories were chosen, each representing a unique type of time series data. By including these datasets, we cover all fifteen types of classification datasets within the UCR repository, namely Device, ECG, EOG, EPG, Hemodynamics, HRM, Image, Motion, Power, Sensor, Simulated, Spectro, Spectrum, Traffic, and Trajectory. These types encompass a diverse range of domains, such as medical signals, motion capture, sensor readings, and simulated environments, among others.
The UCR repository is widely recognized for its extensive collection of time-series datasets, each belonging to a distinct category that reflects different real-world applications. This ensures that our selections are not only classic and representative but also demonstrate the broad applicability and generalizability of our experiments across diverse domains. Our goal was to delve deeper into the nuanced relationships between DA methods and datasets. Selecting fewer, but highly representative datasets, allows us to achieve this objective more effectively.
TABLE II: Selected datasets from the 2018 UCR Time Series Classification Archive
Dataset Type Train Test Class Length
CBF Simulated 30 900 3 128
ECG5000 ECG 500 4500 5 140
FordB Sensor 3636 810 2 500
GunPointAgeSpan Motion 135 316 2 150
ScreenType Device 375 375 3 720
Strawberry Spectro 613 370 2 235
Yoga Image 300 3000 2 426
EOGHorizontalSignal EOG 362 362 12 1250
Fungi HRM 18 186 18 201
GestureMidAirD1 Trajectory 208 130 26 Vary
InsectEPGRegularTrain EPG 62 249 3 601
MelbournePedestrian Traffic 1194 2439 10 24
PigCVP Hemodynamics 104 208 52 2000
PowerCons Power 180 180 2 144
SemgHandMovementCh2 Spectrum 450 450 6 1500
In Table II, the term “Class” refers to the number of distinct categories or labels within the dataset. It reflects the granularity and diversity of events or phenomena that the dataset captures. A higher number of classes generally suggests a more complex classification task. “Length” denotes the duration of each individual time series in the dataset, representing the number of time points or observations for each sequence. This factor is crucial as longer sequences might capture more detailed dynamics but also introduce challenges in computational complexity and potential for overfitting.
All datasets were firstly rescaled to ensure that the minimum value in the training set is -1 and the maximum is 1. This normalization step helps preserve relative values and maintain consistency across different datasets. Additionally, we address missing values, if present, by substituting them with zero, a standard approach that maintains the dimensional integrity of the data while ensuring that missing values do not influence the learning process.
V-B Evaluated Network Models
For our evaluation, we use two deep learning models: the deep convolutional neural network (CNN) model ResNet and the Long Short-Term Memory (LSTM) model. ResNet is particularly well-suited for time series data due to its ability to capture hierarchical features through convolutional layers and its robustness to overfitting via residual connections. LSTM, on the other hand, excels at learning temporal dependencies and capturing long-term trends in sequential data due to its unique memory cell architecture, which can retain information over long periods. Both ResNet and LSTM have been widely adopted in the field of data augmentation for time series classification, serving as classic models with extensive usage and validation.
Many studies have demonstrated their superior performance compared to other models, highlighting their effectiveness in this task. In our survey, over one-third of the papers utilized ResNet, LSTM, or both to evaluate their proposed algorithms, including studies such as [39, 55, 30, 3, 49, 36, 105, 50, 106, 107]. This widespread use underscores the robustness and reliability of these models in benchmarking DA methods for TSC. Our decision to use both ResNet and LSTM stems from their prevalent usage and proven success in many renowned studies. Additionally, the unique characteristics of these models make them particularly suited for time series classification. By employing these models, we aim to provide a reliable evaluation of DA methods and are confident they will demonstrate excellent results with the methods we have collected.
Hyperparameters for the 1D ResNet used in this evaluation were informed by research conducted by Fawaz et al. [3] and Iwana et al. [4], where they noted improvements in time series classification using data augmentation. Similarly, the LSTM configurations were guided by studies [4, 108, 109] that have shown its effectiveness in handling temporal dependencies in time series data.
V-B1 LSTM
As shown in Figure 11, our implementation of the LSTM network is based on the architecture proposed by Karim et al. [109]. It incorporates both LSTM and convolutional layers to capture temporal dependencies and extract meaningful features from the time-series data. Although this setup uses LSTM in a non-standard way, it has been shown to be more effective than traditional configurations [4, 110].
The LSTM component begins by permuting the input shape to align with the multivariate time-step structure required by the LSTM layer. This layer consists of 128 units, followed by a dropout layer with a dropout rate of 0.8, which helps prevent overfitting and enhances generalization. The convolutional component consists of three 1D convolutional layers. The first layer applies 128 filters with a kernel size of 8 and uses the He uniform initializer for weights. The second layer increases the number of filters to 256, with a kernel size of 5. The third layer reduces the filter count to 128, using a kernel size of 3. Each convolutional layer is followed by batch normalization [111] and ReLU activation functions, ensuring stable training and promoting non-linearity. A Global Average Pooling (GAP) layer is used to flatten the output of the convolutional layers, and the results are then concatenated with the output from the LSTM layer. This combination allows the model to leverage both sequential information from the LSTM and spatial features extracted by the convolutional layers. The final output layer uses softmax activation to produce class probabilities, making the network suitable for multi-class time-series classification tasks.
Figure 11: The architectures of the neural networks (LSTM Fully Convolutional Network) used to evaluate the data augmentation methods
V-B2 ResNet
As shown in Figure 12, our implementation of ResNet diverges from the original model [112] by incorporating only three residual blocks with varying filter lengths and omitting max pooling. The first block employs three layers of 64 1D convolutions, the second block contains three layers of 128 1D convolutions, and the third block features three layers of 128 convolutions. Each residual block has three convolutional layers with filter sizes of 8, 5, and 3, respectively. The batch normalization functions [111] and the ReLU activation functions follow each convolution. Additionally, the residual connections merge the input of each residual block with the input of the subsequent block through an addition operation. The final two layers of ResNet comprise a Global Average Pooling (GAP) layer and an output layer with softmax activation.
Figure 12: The architectures of the neural networks (ResNet) used to evaluate the data augmentation methods
All network parameters were initialized using the Glorot uniform method [113]. We trained the network using the Adam optimizer [114], with an initial learning rate of 0.001 and exponential decay rates for the first and second momentum estimates at 0.9 and 0.999, respectively, following Fawaz et al.’s methodology [3]. To maintain consistency in the comparison of data augmentation methods, we ensured that each model underwent the same 10,000 iterations, reducing the learning rate by 0.1 on plateaus of 500 iterations. The number of patterns in the training dataset was quadrupled compared to the original dataset. The choice of 10,000 iterations stemmed from observations that indicated that the training loss for the model on each dataset typically converged before this point.
Moreover, as the number of iterations remained constant, the control experiment without augmentation effectively repeated the same patterns to match the augmented experiments. Regarding hyperparameter selection, we adhered to the best practices documented in the literature as closely as possible, avoiding excessive adjustments to maintain a fair comparison of the data augmentation effects.
V-C Evaluated data augmentation methods
TABLE III: Evaluated Data Augmentation Methods and Their Categories
No. Data Augmentation Method Category Reference
1 None - -
2 Jittering Transformation-Based [6]
3 Rotation Transformation-Based [6]
4 Scaling Transformation-Based [6]
5 Magnitude Warping Transformation-Based [6]
6 Permutation Transformation-Based [6]
7 Random Permutation Transformation-Based [6]
8 Time Warping Transformation-Based [6]
9 Window Slicing Transformation-Based [15]
10 Window Warping Transformation-Based [15]
11 SFCC Transformation-Based [39]
12 SPAWNER Pattern-Based [45]
13 wDBA Pattern-Based [16]
14 RGW Pattern-Based [48]
15 RGWs Pattern-Based [48]
16 DGW Pattern-Based [48]
17 DGWs Pattern-Based [48]
18 DTW-Merge Pattern-Based [55]
19 GAN Generative [66]
20 EMD Decomposition-Based [94]
The data augmentation methods scrutinized in this study are delineated in Table III. The criterion for their selection was predominantly their recurrent citation in the scholarly literature and the accessibility of their open source implementations, which facilitates reproducibility and comparative scrutiny. Amidst a plethora of techniques, some entailing elaborate combinations and complexities, the endeavor was to encapsulate those epitomizing foundational and seminal contributions in the field. This strategic choice ensures the study’s resonance, relevance, and alignment with contemporary scholarly discourse and benchmarks. Each method under evaluation was configured using parameters from their respective seminal papers, thereby preserving the authenticity of their original design and ensuring an equitable evaluation landscape. These methods were evaluated based on their efficacy in enhancing the performance of both the ResNet and LSTM models across various time series datasets.
We excluded Automated Methods due to their complexity and the need for extensive parameter tuning. These methods often integrate multiple algorithms, adding complexity beyond our study’s scope, which focuses on foundational algorithms. The lack of open-source resources also makes reproducibility difficult, potentially compromising reliability. Our aim is to evaluate foundational data augmentation techniques that serve as the building blocks for more advanced models, providing a comprehensive understanding of their capabilities and limitations.
Many generated data augmentation techniques require external training and pre-training, adding significant complexity to the evaluation process. Each dataset often requires tailored training, making it resource-intensive and challenging to maintain consistency across diverse datasets. Evaluating these methods also involves careful hyperparameter tuning and managing potential variability in outcomes, affecting reproducibility and reliability. Due to these challenges and the impracticality of including all methods, we selected GANs as a representative of generated methods, as they offer a manageable balance of complexity within our evaluation framework.
V-D Evaluation Metrics
In addressing the intricate task of evaluating algorithm reliability within the domain of time series analysis, we acknowledge the absence of a singular metric universally applicable across all conceivable applications. This context underpins our selection of evaluation metrics, aiming to bridge the gap between general applicability and the specific demands of our study. This research used a multifaceted evaluation metric system to assess the effectiveness of data augmentation methods. Each metric contributes a unique evaluative perspective, cumulatively offering a holistic insight into each method’s efficacy. The metrics incorporated are as follows:
1. Accuracy (Acc): Accuracy, an intuitive and globally recognized metric, gauges the ratio of correctly predicted instances in all prognostications. We have opted for Acc as a primary metric due to its relevance in scenarios where the correct identification of both true positives and true negatives is of paramount importance. This choice is informed by the nature of our balanced dataset, where equal importance is placed on each class and the distinct types of classification errors are deemed equally significant. In scenarios where false negatives and false positives are critically important, alternative metrics such as the F1-score might be preferred; however, given the balanced nature of our datasets and the equivalence of misclassification impacts, Acc was deemed the most appropriate metric for our analysis of data augmentation in time-series classification tasks.
2. Method Ranking: To facilitate a more nuanced comparative analysis of the data augmentation methods under investigation, we have adopted a ranking paradigm based on accuracy scores. This approach not only simplifies the identification of the most effective strategies but also establishes a clear hierarchy of performance across different datasets. A lower ranking is indicative of superior methodological performance, offering a structured framework to assess the relative success of each data augmentation strategy.
3. Residual Analysis: Residual plots were instrumental in juxtaposing the performance of each method with the baseline (None). Herein, a residual denotes the deviation between the actual and forecast accuracy. Positive residuals signify outperformance over the baseline, whereas negative residuals indicate underperformance. These plots are particularly suited to our study as they provide a visual conduit for assessing the relative performance increments or decrements attributable to each data augmentation strategy. By visualizing these differences, we can identify patterns and trends in the data that may not be apparent through traditional accuracy metrics alone.
This multimetric evaluation approach underpins the study’s commitment to a nuanced, rigorous, and exhaustive assessment, fortifying the conclusions regarding the effectiveness of each data augmentation strategy.
V-E Research Questions and Objectives
Navigating the complexities of time-series data augmentation necessitates a profound understanding of the subtleties and ramifications of various methods. This study is predicated on the pursuit of clarity in this domain, aiming to elucidate the efficacy and implications of various data augmentation strategies through the following research questions.
RQ 1: What is the impact of data augmentation on model performance?
This pivotal question seeks to explore the influence of data augmentation on model efficacy, probing the disparities in outcomes engendered by different techniques. It bifurcates into more specific inquiries:
(1) RQ 1.1: What is the extent of the universality of data augmentation strategies?
This inquiry is dedicated to discerning the consistency of data augmentation’s effectiveness across various contexts, highlighting any discrepancies and underlying rationales for performance enhancement or degradation.
(2) RQ 1.2: Which data augmentation strategy is the most effective for TSC?
This probe focuses on identifying the preeminent strategy among the evaluated techniques, offering critical insights to practitioners in the time-series domain.
RQ 2: How do intrinsic dataset characteristics influence the effectiveness of data augmentation and subsequent recommendations?
This exploration delves into the nuanced interplay between dataset attributes and augmentation efficacy, aiming to extrapolate actionable insights for optimal method application based on specific dataset traits. The objective is to equip practitioners with informed and customized augmentation strategies, thereby maximizing model performance and circumventing the pitfalls of indiscriminate augmentation application.
The ensuing chapter, “Results Analysis”, will systematically address these queries, contributing to a robust, comprehensive understanding of time-series data augmentation and its diverse implications in machine learning.
VI Results Analysis
In this section, we undertake a methodical examination of the research questions posited at the outset. Our first point of inquiry assesses the universality and optimality of various data augmentation techniques (RQ 1). Subsequently, we explore the nuanced interdependencies between the attributes of the dataset and the efficacy of these augmentation methods, providing customized guidance for practitioners (RQ 2). This discourse contributes a lucid and empirically grounded perspective on the dynamic between data augmentation strategies and time series classification.
VI-A Answer to RQ 1.1: How universal are data augmentation techniques?